### v1

# Dynamic read & write optimizaton with TurtleKV

Abstract.
[…]
1. Introduction
Key-value storage is a foundational building block upon which many data-management applications are built. Though key-value storage’s dictionary-based interface flexibly supports a diverse range of applications, concrete implementations must trade off between the cost of reads, writes, and memory usage. Prior work hypothesizes that only two of these dimensions can be optimized at the cost of the third (Brodal and Fagerberg, 2003; Athanassoulis et al., 2016). The result is a proliferation of implementations tuned for specific workload characteristics (Meta Platforms, Inc., 2022; Conway et al., 2020; Conway et al., 2023; Apache, 2025; Google, [n.d.]). This growing diversity complicates database management: organizations must often deploy multiple databases to support different workloads and rely on expensive Extract-Transform-Load (ETL) pipelines to move data (Databricks, [n.d.]).
Figure 1. Today’s key-value engines (left) trade reads directly against writes, then compensate for lost performance by spending memory. In contrast, TurtleKV(right) uses a balanced on-disk data structure. It dynamically trades off memory for increased write or read performance.
With respect to the tradeoffs, state-of-the art databases limit memory usage and choose between reads or writes for the remaining dimension they wish to optimize (left diagram in Figure 1). They use on-disk structures biased for performance in this remaining dimension and use other mechanisms to recover lost performance in the unoptimized dimension. For example, SplinterDB (Conway et al., 2020; Conway et al., 2023) uses a on-disk data structure, STBϵ-trees, biased toward writes and uses a limited amount of memory for quotient filters/maplets to improve read performance. TreeLine (Yu et al., 2022) uses update-in-place grouping mechanisms and spends memory on histograms to improve reads. Industry approaches, such as RocksDB (Meta Platforms, Inc., 2022), optimize for writes via LSM-trees and use Bloom filters to improve read performance. These approaches are static. They find a single read/write optimization point and permanently change the on-disk data structure, creating a residual memory cost for the entire lifetime of the data.
Though powerful, the above approaches curtail the maximum achievable performance in the unoptimized dimension. As a consequence, supporting load-then-read workloads with high performance requires either spending significant I/O changing the on-disk format or using an expensive ETL pipeline to move data between different databases.
We present TurtleKV, a key-value engine that uses an unbiased on-disk structure suited for reads and writes. It includes a tuning knob that dynamically controls the allocation of memory resources on demand within a fixed budget to improve read or write performance as needed (right diagram in Figure 1). The knob allows memory usage to be re-allocated to the alternate dimension (e.g., reads) when high performance in currently-chosen dimension is no longer needed (e.g., writes). TurtleKV is made possible by the following two observations. First, though the trade-off between memory, reads and writes may be fundamental, the choice of which dimensions to trade-off against are not. Second, many common workloads do not need high write and read performance simultaneously, so memory can be fungibly exchanged between these two dimensions as needed. For workloads that require high read-and-write performance simultaneously, TurtleKV allows human operators to adjust the tuning knob in a weighted way to reflect their operational goals.
To support TurtleKV, we introduce a new external memory dictionary data structure, the TurtleTree, which improves read performance and bounds worst case write I/O behavior. TurtleTrees are a fusion of B+-trees and LSM-trees, leveraging the strengths of both to achieve an unbiased state, balanced between read and write performance. B+-trees, which offer high read performance, provide the high-level structure with data organized in tree-node pages. LSM-trees, which offer higher write performance, are used within the tree-nodes to buffer updates. By carefully managing the TurtleTree structure, we can control the rate at which updates flow down the tree, bounding the number of I/Os required for reads and writes. This property results in a balanced on-disk structure that can be dynamically tuned for greater read or write performance.
We identify TurtleTrees as belonging to a previously unidentified category of data structures, which we name Bϵ+-trees. Such Bϵ+-trees differ from Bϵ-trees and other buffered B-trees by adding a level of indirection between tree-node pages and pages containing per-node buffered updates. SplinterDB’s size-tiered Bϵ-trees (STBϵ-trees) also belong to the Bϵ+-tree family. TurtleTrees differ from STBϵ-trees in that they use level-tiered LSM-trees instead of size-tiered ones. This change shifts the focus from write performance to more unbiased point between reads and writes, while also allowing for more fine-grained control over write amplification. In contrast to traditional level-tiered LSM-trees, TurtleTree’s design limits write I/O costs by bounding the number of pages flushed to disk at any time. Furthermore, TurtleTrees and other data structures in the Bϵ+-tree family are complementary to existing read/memory trade-off mechanisms such as page caching and AMQ filtering because of the separation of data (leaf pages and buffer segments) from metadata (interior nodes and any filters).
Checkpoint distance ($\chi$), which is TurtleKV’s dynamic tuning knob, controls the portion of TurtleKV’s total memory devoted to improving write performance and can achieve the performance of write-optimized, size-tiered approaches on demand. Concretely, this knob sets the number of in-memory batch updates that can be performed on a TurtleTree before updates are written to disk and combines synergistically with TurtleTree ’s control over flushed pages. A higher value allows more data to be merged in memory, reducing write amplification and thus increasing write performance. To dynamically improve read performance when under cache pressure, TurtleKV can choose a smaller value, so that memory that is unused for writes maybe (re-)allocated to cache more data pages as needed. The key benefit of this knob is that unlike existing systems that permanently spend memory by creating new structures such as filters, tuning decisions are not anchored to on-disk structures at the point data is written. TurtleKV can dynamically optimize reads or writes by simply adjusting the allocation of memory.
[…]
Key contributions:
  1. 1)
    An exploration of the benefit of trading off memory dynamically for increased write or read performance. The design of an architecture that enables high dynamic tuning range by starting with a unbiased on-disk data structure and dynamically trading memory for performance.
  2. 2)
    TurtleTrees, a novel on-disk data structure that merges the benefits of read-optimized B+-trees with write-optimized Level-tiered LSM-trees. The identification and formal definition of an important class of data structures to which TurtleTrees belong, Bϵ+-trees
  3. 3)
    The concept of checkpoint distance ($\chi$), which allows more batched updates to reside in memory before being written to disk. This limits write amplification as previous batches can expire in memory before being written.
  4. 4)
    TurtleKV, a key/value storage engine that uses TurtleTrees and checkpoint distance to achieve high levels of performance on modern SSDs.
2. Overview
Brodal and Fagerberg’s Lower Bounds (Brodal and Fagerberg, 2003) and the RUM Conjecture (Athanassoulis et al., 2016) show that key/value stores are constrained by a three-way trade-off between main memory usage, read operations, and writes. This means that storage engines have at most two degrees of freedom when selecting a tuning point for a given application: for example, committing to a cost function for reads and memory forces a minimum cost for writes.
A key insight of our work is that, while this trade-off is fundamental, the mechanisms in common use by today’s storage engines to navigate the RUM space are not. This means it is theoretically possible to devise a dictionary type affording mechanisms which trade, for example, writes against memory independent of reads, and reads against memory independent of writes. The goal of our work is to demonstrate mechanisms which do this in a way that avoids coding a given tuning point into the structure of stored data. This allows tuning points to be chosen dynamically, by re-allocating memory in different ways to optimize for the workload at hand, expanding the range of applications which can be supported by a single general-purpose key/value store.
In this section we give a brief review of current practice with respect to trade-off tuning in general purpose key/value storage engines, grouped by tuning goal: read-optimized, write-optimized, and hybrid.. We then present the key ideas of this work: Bϵ+-trees, TurtleTrees, Write Optimization via Checkpoint Distance, and On-Demand Optimization.
2.1. Current Approaches to Trade-off Tuning
Here we describe and critique the major strategies presently in use to navigate the RUM trade-off space based on data structure type.
Figure 2. Bϵ-tree $\rightarrow$ Bϵ+-tree structural transformation
Read-Optimized (B-trees). B-trees (Comer, 1979) provide optimal (Yi, 2012) read performance and can utilize memory to reduce the number of reads required to process queries by caching pages containing tree-nodes. B+-trees, the most commonly implemented B-tree variant, exploit this mechanism by placing key/value data in the leaves, allowing all but the bottom layer of the tree to be cached in main memory so that queries can be answered with a single read.
State-of-the-art approaches to reducing the write cost of updates in B-trees employ the use of a write-ahead log (WAL) to ensure durability of updates, plus a scheme for applying updates to the tree in memory. We refer to these as Log Structured B-trees. Their goal is to defer rewriting pages containing updated keys, in hopes of collecting enough updates per page to amortize I/O. Bw-trees (Levandoski et al., 2013a) do this through the use of delta records applied in-memory to the main tree using lock-/latch-free techniques. The closely related Bf-tree (Hao and Chandramouli, 2024) use an alternate strategy of caching and updating sub-page slices (mini-pages) instead full pages, which has additional benefits in terms of lowering the memory cost of caching ”hot” records.
Bw-trees and Bf-trees achieve good write efficiency either when memory is plentiful or when updates are clustered in a few sub-ranges of the key space. In the general case, however, Log Structured B-trees force a sub-optimal trade-off between memory and write amplification, especially as the total data size $N$ grows with respect to available memory $M$ (Idreos et al., 2019).
Write-Optimized (LSM-trees). Log-Structured Merge (LSM)-trees (O’Neil et al., 1996) are a popular write-optimized alternative to B-trees. By deferring the merge of new updates with existing data, they maximize the amount of effective work achieved by each write operation. Unfortunately, this technique forces read costs (and space requirements) to increase, since it introduces the possibility that updates for a given key may be stored in multiple locations. Page caching can be used to trade memory against reads, but this mechanism is less effective for LSM-trees than B-trees because there is potentially much more data to cache.
In addition to caching, two primary tuning mechanisms are employed by LSM-Tree systems: compaction policy and approximate memory query (AMQ) filtering. Compaction policies divide into greedy (level-tiering) and lazy (size-tiering). In a level-tiered LSM-Tree, write cost is increased by repeatedly merging data at the same level (up to $F$ times) to reduce future read cost. In a size-tiered LSM-Tree, a tunable number $T$ of key duplicates are allowed to co-exist at a given level before their eventual merge into a larger level. Both compaction policies directly trade read cost against writes by altering the invariants of stored data.
Approximate member query (AMQ) filters, e.g. Bloom Filters (Bloom, 1970), allow memory to be spent to reduce the read cost of point queries by having a dedicated filter (or part of a shared filter (Conway et al., 2023)) devoted to each location where a key might reside. AMQ filters offer a tuning parameter $\alpha$, the false-positive rate, which trades size against accuracy. [^1] A filter must always answer accurately if a query key is not in its set, but it may give a false positive result at rate $\alpha$. The smaller $\alpha$ gets, the more bits per key are required to store the filter. To answer a point query for key $k$, the filters for all locations $S$ where $k$ might reside are consulted before performing any I/O. This rules out, with probability $\prod_{i\in S}{\left(1-\alpha_{i}\right)}$, locations where the query will fail, preventing excess reads.
If filters are used to maintain a target read cost in a size-tiered LSM-tree (with $T$ tiers per level), the total size of the filters must scale in proportion to $T$, since the total number of false-positives for a query is the sum of false-positives for all filters that pertain to a given key (Dayan et al., 2018). Therefore the $T$ parameter in a size-tiered LSM, together with $\alpha$, can be thought of as representing a trade-off between memory and write amplification. The problem with this mechanism is that the future read amplification one trades to achieve better write performance is reflected in the on-disk structure, and thus the memory cost for filters to compensate must be paid during the entire readable lifetime of write-optimized updates. Note that this is true even if the filter type (e.g. Quotient Maplets (Conway et al., 2023)) supports cache-friendly paging of partial filters into main memory, since the filter bit rate must nevertheless be increased to maintain a target query cost, regardless of how much of a filter is cached. In essence, write-optimization in the past generates a read/memory trade-off penalty in the future in such designs. We refer to this effect as the residual write-optimization penalty.
Hybrid-Optimized (Bϵ-trees). Bϵ-trees (Bender et al., 2015; Brodal and Fagerberg, 2003) are a tunable external memory dictionary, able to trade reads against writes. The basic B-tree structure is used as the foundation for a Bϵ-tree. The $\epsilon$ parameter varies the portion of a node page used for pivots, which optimize for reading, and update buffering, which optimizes for writing. At either extreme of their tuning range $\epsilon\in{0..1}$, Bϵ-trees simply degrade into the equivalent of a B-tree or LSM-tree. At $\epsilon=\frac{1}{2}$, the branching factor is $\sqrt{B}$, making tree height twice that of an equivalent B-tree. In this configuration, updates can be flushed down one level at a time from the root in batches of size $\Theta(\sqrt{B})$, dramatically lowering the worst-case write cost and providing a good balanced tuning point.
One disadvantage of traditional Bϵ-trees, where updates are buffered inline within the node page (Figure 2, 3 ), is that, for any given value of $\epsilon$, increasing the page size $B$ to optimize writes increases the I/O and/or memory cost of reading/caching a node. Similarly, decreasing $\epsilon$ to decrease the branching factor and increase the buffer size increases the height of the tree, which raises the cost of reading/caching a node-path to any given key.
Size-Tiered Bϵ-trees, used by SplinterDB (Conway et al., 2020; Conway et al., 2023), are a recent re-imagining of the Bϵ-tree designed to close the ”bandwidth gap” between the capabilities of fast NVMe SSDs and traditional key/value storage engine designs. The main structural modification in STBϵ-trees is to add level of indirection between node pages and buffered data (Figure 2, 4 ). This change is analogous to the transformation between B-trees and B+-trees, and gives SplinterDB similar advantages in terms of being able to cache interior nodes for reading.
SplinterDB adopts the strategy of selecting a single tuning point which attempts to offer excellent read and write efficiency by structuring per-node buffer data pages as a single-level, size-tiered LSM-tree-like structure, where each sorted run (or branch) is stored as a statically built B+-tree. Quotient Maplets, a variant of Quotient Filters, are used in this design to route point queries to the branches/tiers which may contain a given key.
While benchmark performance data (YCSB) shows this to be a good compromise for point-query workloads on modern hardware, it nonetheless represents a trade-off between false positives leading to increased read amplification on the one hand, and the memory cost of maintaining maplets with sufficient bits per key to compensate for the number of possible redundant copies of a key stored at multiple tiers. The downside of this approach is felt most acutely in short (¡100 key) range scans, which must merge/compact a potentially large number of tiers to determine sequential keys.
2.2. Towards Dynamic Tunability: Key Ideas
In this work, we propose a dynamic, unbiased approach that achieves desired read and write performance on demand while avoiding the costly re-tuning or data movement required in existing, statically-tuned systems.
Key Idea 0: Bϵ+-trees. A baseline requirement for dynamic optimization is an unbiased data structure with sufficient flexibility to achieve both high write and read performance. To achieve this property, we identify and build upon what we consider to be the core insight behind SplinterDB’s STBϵ-trees: moving buffered updates in a Bϵ-tree out of the node pages and into pages referenced from nodes (Figure 2, 3 $\Rightarrow$ 4 ). Following the observation that this transformation is quite similar to the difference between B-trees and B+-trees (Figure 2, 1 $\Rightarrow$ 2 ), we coin the term Bϵ+-tree to describe this new class of dictionaries.
Moving update data out of node pages improves flexibility by offering potential advantages for both read- and write-optimization in Bϵ+-trees. Read efficiency benefits from smaller nodes being more easily cached; likewise, queries benefit from reduced tree height as a result of a higher branching factor, since the portion of a node devoted to pivots can be larger without sacrificing buffer capacity. In terms of write efficiency, being able to buffer more data per node allows updates to be moved down the tree in larger batches, allowing more work to be achieved per write during flushes. The identification of Bϵ+-trees opens the door for research into a set of important design dimensions around how the data within a node buffer is structured. For example, instead of the STBϵ-tree’s flat structure of references to static B+-trees with size-tiered compaction policy, adopting a more greedy (level-tiered) compaction policy across levels within a node shifts away from a write-optimized focus towards a more unbiased one. A key contribution of this work is to begin exploring some of this space of possibilities to better understand their potential.
Key Idea 1: TurtleTrees. With Bϵ+-trees as a foundation, we introduce a novel data structure that leverages an unbiased, level-tiered node buffer structure and flush policy to enable dynamic read and write optimization. We name this new data structure the TurtleTree, shown in Figure 3. TurtleTrees are an instantiation of Bϵ+-trees in which buffer c -referenced data segments d are stored as flat leaf-like pages organized into a small, fixed capacity level-tiered LSM-tree per node a . Unlike size-tiered Bϵ+-trees, which are biased towards improving write performance at the cost of reads or log structured B-trees, which are biased towards improving reads over writes, the level-tiered TurtleTree sits at a middle point between these two poles. Like LSM-trees and other Bϵ-tree variants, TurtleTrees permit updates for a given key to be stored in multiple places (though keys must be unique within a given buffer level). By utilizing a level-tiered compaction rather than size-tiered, the number of potential locations is reduced, accelerating reads. TurtleTrees can also still use the various memory/read trade-off mechanisms discussed earlier, such as page caching and AMQ filtering, to further reduce read cost. In fact, we will see in this work that TurtleTrees can achieve read performance that is on par with or exceeds current state of the art read-optimized key/value stores.
Figure 3. The TurtleTree data structure.
Setting aside reads, the question remains as to how to restore the write efficiency we lost by changing the intra-node compaction policy to level-tiered. We present a mechanism which achieves this in the next key idea (2). In preparation, we highlight a unique characteristic of TurtleTrees. By building a level-tiered structure within a Bϵ+-tree, we can bound the I/O cost of writing batches of modified data pages from the TurtleTree in-memory to disk, avoiding the write amplification penalty of traditional level-tiered LSM-trees. In fact, by setting the number of data pages d referenced by the buffer c roughly equal to the number of pivots b  [^2] The diagram in Figure 3 is conceptual; the exact ratios of pivots to buffer segments, and their relative sizes, are not accurate/to-scale., we guarantee that for each batch of pages inserted in the TurtleTree, we can flush an equal-sized batch to a lower level, bounding the growth of write amplification. This fine-grained control over how writes propagate down a TurtleTree allows us to choose exactly how many pages to flush and when to flush them.
Key Idea 2: Write Optimization via Checkpoint Distance. Given the unbiased TurtleTree structure, we can now introduce checkpoint distance, a tuning knob that improves write performance in exchange for memory without affecting future performance via residual optimization penalties. When receiving a batch of updates, TurtleKV applies the batch to a TurtleTree in-memory and can immediately write the modified pages to disk, creating a checkpoint. However, it is also possible to defer this write, buffering the changes from multiple batches within the TurtleTree in-memory before creating the corresponding checkpoint on disk. We refer to the number of batches $\chi$ buffered in-memory as the checkpoint distance. By modifying $\chi$ at runtime, it is possible to trade write performance for memory in TurtleKV; specifically, the number of writes required per key update is reduced by $O\left(\log{\chi}\right)$! This reduction is independent of the total size $N$ of stored data, making it a viable tuning knob regardless of the on-disk size or structure.
To understand why this is the case, consider the lifetime of one of the buffer data pages d shown in Figure 3. As new batches are applied to a node, the merging/compaction process causes data pages at higher levels to be frequently replaced with merged data and written to lower levels. This means that the closer a buffer segment’s level is to the root, the shorter its effective lifetime. By buffering batches in the TurtleTree, these short-lived data pages naturally expire and no longer need to be written to disk. Leveraging this property, we arrive at a key insight: by spending memory to defer writing these short-lived pages, we can avoid unnecessary writes and reduce write amplification without changing the stored structure of the eventual checkpoint.
Prior work such as SplinterDB (Conway et al., 2020; Conway et al., 2023) or TreeLine (Yu et al., 2022) permanently spend memory by storing additional filters or histograms to index into their data pages, incurring a residual optimization penalty each time they want to optimize performance. In contrast, checkpoint distance leverages the TurtleTree structure to improve write performance with no additional memory cost beyond the tree itself and its internal metadata. This means that checkpoint distance can be adjusted up or down to control write amplification without incurring any residual optimization penalty.
Key Idea 3: On-Demand Optimization. Instead of one-time, static optimization, TurtleKV leverages the TurtleTree structure to optimize both reads and writes on-demand, creating a general-purpose KV engine that negates the need for expensive data movement present in current Extract-Transform-Load (ETL) pipelines. TurtleTrees offer a tunable external memory dictionary which can navigate the RUM trilemma space using only read vs memory and write vs memory tuning mechanisms. We take advantage of this property by using checkpoint distance as a tuning knob to dynamically determine how much of the total available memory to spend on optimizing writes or reads. When the checkpoint distance is larger, more memory is spent buffering batches within TurtleTrees, optimizing writes, while when the checkpoint distance is smaller, more memory is dedicated in the page cache for data pages, optimizing reads. Since the tuning mechanism trades reads and writes against memory rather than reads vs writes directly, adjusting the memory allocation does not affect the underlying data structure. Thus, TurtleKV can dynamically adjust the memory allocation without permanently affecting either read or write performance, allowing for on-demand optimization of reads or writes depending on application requirements.
In systems whose tuning mechanisms trade reads against writes directly, such as compaction policy in LSM-tree storage engines, the cost of re-tuning is high because it involves changing the stored structure of data to target a new trade-off point. Historically, applications have implemented this kind of ”tuning by moving” strategy through the use of multiple data stores and Extract-Transform-Load (ETL) pipelines, which transfer data between storage systems to optimize it for various application requirements. The fact that ETL and multi-stores have been quite popular despite high cost and increased system complexity speaks to the need for data management systems which can re-tune on demand to support changing workloads. In contrast to expensive data movement, TurtleKV can simply spend memory whenever it is advantageous to do so to accelerate read- or write-heavy workloads: pay now, not later.
3. TurtleTrees & Checkpoint Distance
[…]
Compatibility with other mechanisms. We focus on $\chi$ as the main parameter to improve TurtleTrees’ performance. TurtleTrees are also compatible with other mechanisms to increase read performance: increasing page-cache size and read-filter bitrate. In this work, we assume static sizes for these additional parameters to isolate the effect of TurtleTrees and $\chi$.
3.1. TurtleTree Structure
TurtleTrees are Bϵ+-trees ($\S$2.2) whose per-node update buffer is organized as a fixed-capacity, level-tiered LSM-tree with fanout F=2F=2.333That is, level to level growth factor; not to be confused with the branching factor of the node, or number of pivots, which we denote as $\rho$. The structure of TurtleTree node and leaf pages is shown in Figure 4.
Leaf pages contain a sorted list of key/value pairs. As is typical of B+-trees, leaves are sized several orders of magnitude larger than interior nodes. [^4] A typical configuration might be 4KB node pages and 16MB leaf pages. To ensure fast in-memory queries (both point and range/scan), TurtleKV leaf pages contain a cache-oblivious, byte-addressed binary trie index. This index contains every 16th or 32th key from the leaf, depending on available space. The trie is made cache-oblivious by laying out nodes in Van Emde Boas order (Bender et al., 2007) to maximize locality of reference along any given query path.
Each (interior) node is split into roughly equal halves, with one half devoted to pivots (dark gray) and the other to the update buffer (yellow). Within a node page, the update buffer region contains pointers to segments (the blue boxes with rounded sides), which are stored as leaf pages with the addition of tombstone records for deleted keys. The segments in a buffer are stratified into levels of exponentially increasing size. Keys are required to be unique and sorted within a level. The first level contains at most a single segment, the second level at most two, the third level four, then eight, etc. All levels start out vacant and can be in either a vacant or occupied state. For a given number of pivots $\rho$ in a node, the number of levels is set to $\lceil\log_{2}{\rho}\rceil$, and the total number of segments in all levels of the buffer bounded to at most $\rho-1$.
Figure 4. Details of the internal structure of TurtleKV’s TurtleTree implementation.
Node buffers maintain certain metadata to speed up various operations. For each segment, we store a bit set activePivots of the pivots addressed by at least one active key in the segment, [^5] We say that a pivot is addressed by a key if the pivot’s key range contains that key. and a compressed sparse array flushedPivots of key index upper bound values for each pivot to which keys from the segment have been flushed. When keys are removed from a level as part of a flush operation, the segment pages are not modified; instead, the flushed upper bound array is updated for that pivot so it points to the segment’s first non-flushed key in the flushed pivot key range. When all keys addressed to a given pivot have been flushed, the active bit is cleared for that pivot, and the flushed upper bound is reset to zero. If this causes the active set to become empty, the segment is removed from the node, possibly allowing its storage to be garbage-collected. [^6] Operations such as node splits during a batch update or taking a snapshot of an existing TurtleTree may introduce additional references to a given page.
Node pages also maintain a count of the number of bytes of buffered updates in the key range of each pivot. This is used to decide when and how to perform flush operations. The default policy is to flush one leaf-sized batch after each batch insert, to the pivot with the highest pendingBytes count, if possible.
3.2. TurtleTree Batch Updates
TurtleTrees are designed to support efficient batched updates; individual key modifications can be logged/buffered outside the tree and applied as a batch once enough have accumulated. As such, TurtleTrees do not directly implement single-key Put or Delete.
When the tree is empty, or consists of a single leaf, most operations are trivial. Batch update on a leaf consists of doing a merge/compact on the new and old data; if the result is small enough to fit in a leaf page, then the tree remains a single leaf. If not, then a new node is created to parent the two leaves. The root node is the only node in the tree allowed to have an arbitrarily small number of child pivots, so long as it is at least 2.
Figure 5. Node-Buffer Insertion of Batch Updates in TurtleTrees.
To perform a batch update on a node, we first insert the incoming batch into the node’s buffer. To illustrate how this works, we work through a simple example, shown in Figure 5. In this example, we process four batches, containing keys $\langle 1,7,10\rangle$, $\langle 0,4,5\rangle$, $\langle 2,8,11\rangle$, and $\langle 3,6,9\rangle$ respectively. The node’s buffer is initially in an empty state, i.e. all levels are vacant.
When the first batch $\langle 1,7,10\rangle$ is inserted, we look at the first level of the buffer to see if it is vacant or occupied. It is vacant, so we can simply insert the batch as a new segment page a. At the second batch, $\langle 0,4,5\rangle$, we again check the first level, but this time it is occupied. Thus, we must perform a merge of the keys in segment a and batch 2 to produce the sequence $\langle 0,1,4,5,7,10\rangle$, which is split into two segments b $\langle 0,1,4\rangle$ and c $\langle 5,7,10\rangle$. The third batch $\langle 2,8,11\rangle$ becomes a new segment d occupying the first level, just as the first batch. When the fourth batch $\langle 3,6,9\rangle$ is inserted, it must be merged with both levels one and two, since both are occupied. This cascade of merges produces the final sorted key sequence $\langle 0,1,2,3,4,5,6,7,8,9,10,11\rangle$, resulting in a new level comprised of segments e $\langle 0,1,2\rangle$, f $\langle 3,4,5\rangle$, g $\langle 6,7,8\rangle$, and h $\langle 9,10,11\rangle$.
After inserting a new batch into the buffer, if there are at least $L$ (the size of a leaf page) buffered bytes of updates addressed to one of the pivots, then we perform a flush. This involves picking the pivot with the most buffered data, and removing $L$ bytes of updates from the buffer in the target key range. Extracting a batch to flush may involve an in-memory merge/compaction, as the data may be spread across multiple levels. In the prior example, if we perform a flush after batch 3, supposing the target subtree’s key range to be $2..6$, we would need to merge data from all three segments (b, c, d) to produce the batch $\langle 2,4,5\rangle$. As mentioned in $\S$3.1, durable data in update buffer segments is removed by modifying the in-node active key range metadata. This reduces write amplification by avoiding the need to update the segment page(s).
Operation TurtleTree B+-Tree $x$ LSM (level-tiered) LSM (size-tiered) Bϵ-Tree STBϵ-Tree
Put $\frac{1}{B}\cdot\log_{2}{\frac{N}{\chi}}$ $\log_{B}{N}$ $\frac{F}{B}\cdot\log_{F}{N}$ $\frac{1}{B}\cdot\log_{T}{N}$ $\frac{1}{B^{1-\epsilon}}\cdot\log_{B^{\epsilon}}{N}$ $\frac{1}{B}\cdot\log_{F}{N}$
Get (DAM) $\log_{2}{N}$ $\log_{B}{N}$ $\log_{F}{N}\cdot\log_{B}{N}$ $T\cdot\log_{T}{N}\cdot\log_{B}{N}$ $\frac{1}{\epsilon}\cdot\log_{B^{\epsilon}}{N}$ $F\cdot\log_{B}{N}$
Get (PDAM) $\log_{B}{N}$ $\log_{B}{N}$ $\log_{B}{N}$ $T\cdot\log_{B}{N}$ $\frac{1}{\epsilon}\cdot\log_{B^{\epsilon}}{N}$ $F\cdot\log_{B}{N}$
Table 1. Per-key, asymptotic costs (Big-$O$) for external memory dictionaries. N = number of inserts, B = I/O transfer (block) size, F = level growth factor, T = tiers per level, $0\leq\epsilon\leq 1$ controls the fraction of a node allocated for pivots ($B^{\epsilon}$) vs buffer ($B^{1-\epsilon}$). Point-Query (Get) costs are given both for Disk Access Model (DAM) and Parallel Disk Access Model (PDAM). We omit the $L$ (leaf size) parameter and assume $\rho=O(B)$, to simplify the expressions and to ease comparison across data structures. Scans (length $k$) cost a Get plus $\frac{k}{B}$ I/Os for all cases and models.
After a batch has been recursively flushed to a sub-tree, a split or join of the corresponding pivot may be required. All non-root nodes must be kept at least half-full to remain viable. In addition, we maintain an invariant that the total amount of buffered data per node never exceeds $L\cdot(\rho-1).$ If the buffered data does ever exceed this amount, it will be because a new batch of size $L$ was just inserted, resulting in there being at least $L\cdot\rho$ total bytes. In this case, we will always have enough to flush a batch to some sub-tree to restore the invariant. Node joins are simpler than splits, since we can assume the invariant to hold for each node being joined. Since the key-range of two sibling nodes is guaranteed not to overlap, the invariant is preserved by simply concatenating the contents of nodes to join them. The same is not true, in general, of splits. If after splitting a node as evenly as possible, we find that the update buffer size invariant is violated for one of the new smaller nodes, we perform another flush on that node to restore it. Note that we can be sure that if this case arises, it will only be for one side, not both; otherwise, the invariant must have been false before the split. A quotient filter per leaf page is built and flushed when batch updates are finally written to storage.
3.3. Checkpoint Distance
Figure 6. The effect of checkpoint distance ($\chi$) on write amplification. (Pages shown are from the example in Figure 5.)
The $\chi$ parameter controls how frequently newly created pages (nodes, segments, and leaves) are written from memory to storage. At $\chi=1$, every time we perform a batch update on the tree, all modified pages must be written back out. At $\chi=2$, we defer the writing of new/modified pages until two batches have been applied, etc. It may not be immediately obvious to the reader that $\chi$ should have any effect on the write cost of key updates. To see that it does, let us revisit the batch insertion example illustrated by Figure 5, considering the span of time (as measured by batch number) that each page is live, shown in Figure 6. The pages which appear in durable storage at a given time step are those whose lifespan overlaps a given column in the middle section (Lifetime) of the diagram.
Observe that when $\chi=2$, the state of the tree is only externalized at time steps 2 and 4, so that the segments (a, d) which are live only for steps 1 and 3 are never written. Also observe that when $\chi=2$, keys 1, 7, and 10 first appear in stored pages at level two in segments b and c; i.e., they appear to an external observer (who can only look at the externalized or durable states of the tree) to have skipped the first level altogether. If one considers batch 3, the effect is even more pronounced: the first place keys 2, 8, and 11 appear when $\chi=2$ is at the third level, in segments e, g, and h. If $\chi$ is doubled again to 4, then all keys in the example appear to skip the first two levels.
3.4. Cost Analysis & Limitations
Table 1 compares TurtleTrees to other data structures. We see that TurtleTree’s update put (update) operations benefit greatly from checkpoint distance due to the $\chi$ parameter in the denominator. Gets under different memory models perform similarly to other data structures, but are greatly increased by read caching (not shown in the table). We next describe how the table values are derived.
Individual key updates travel from root to leaf of the tree as batch updates and flushes take place. Since there are $O\left(\log_{2}{\rho}\right)$ levels within the buffer of each node, and the tree height is $O\left(\log_{\rho}{\left(N/L\right)}\right)$, each key update is written $O\left(\log_{2}{\rho}\cdot\log_{\rho}{\left(N/L\right)}\right)=O\left(log_{2}{\left(N/L\right)}\right)$ times, regardless of the node page size or tree branching factor $\rho$. Note that in order to correctly analyze the cost of a key update, we need only consider its initial journey from root-to-leaf, since each time a key’s page is re-written after it reaches the bottom of the tree, we can amortize that cost against the incoming batch, which represents an equal number/size of keys ending their root-to-leaf journey. If we define the transfer size $B$ in units equal to a single key update, the I/O cost of a key update is
$O\left(\frac{1}{B}\cdot\log_{2}{\frac{N}{L}}\right)$
Point queries (Get) differ in their I/O cost depending on whether or not the key is found. Here we give a worst-case bound for the general case; in practice this can be dramatically improved through the addition of per-leaf filters. The cost analysis here also depends on the specific I/O model used; if we are counting total page loads, as in the Disk Access Model (DAM), then a single-key (point) query must load all node pages and buffer segments along a single root-to-leaf path: $O\left(\log_{\rho}{\frac{N}{L}}+\frac{L}{B}\cdot\log_{2}{\frac{N}{L}}\right)$ or just $O\left(\frac{L}{B}\cdot\log_{2}{\frac{N}{L}}\right)$.
However, because “latency lags bandwidth” (Patterson, 2004), it is also reasonable to consider I/O models where latency and parallelism are explicitly accounted for (Bender et al., 2021, 2019; Aggarwal and Vitter, 1988), e.g. the Parallel Disk Access Model (PDAM). Under the PDAM model, Turtle Trees only incur a cost of $O\left(\log_{\rho}{\frac{N}{L}}\right)$ parallel I/Os per point query: one per node on the key’s root-to-leaf path. This makes point queries on Turtle Trees on-par with (non-buffered) B-trees.
Like other B-tree variants, range queries on Turtle Trees cost a point query (to find the starting key) plus $O\left(\frac{k}{B}\right)$ IOPs. There is also an in-memory, per-key cost of $\log_{2}{(\log_{2}{(N/L)})}$ to merge sorted runs, which is the same as a level-tiered LSM-tree. This per-key cost expression is derived via the use of a binary min-heap to select which of the $\log_{2}{N/L}$ levels of the tree contains the least-ordered key.
How $\chi$ effects cost analysis. We claim that, in general, the effect of $\chi$ is that keys skip the first $\log_{2}{\chi}$ buffer levels of the tree, which reduces the per-key write cost to:
$O\left(\frac{1}{B}\cdot\left(\log_{2}{\frac{N}{L}}-\log_{2}{\chi}\right)\right)=O\left(\frac{1}{B}\cdot\log_{2}{\frac{N}{\chi L}}\right)$
Limitations. The key limitation of TurtleTrees and $\chi$ is memory usage to support $\chi$ instead of read caching. But, $\chi$ can be tuned dynamically during runtime so $\chi$ need only be increased when needed. In return, TurtleTrees and $\chi$ pay less read-filter cost for the lifetime of data.
4. From TurtleTrees to TurtleKV
Figure 7 shows TurtleKVs high-level architecture. It is similar to systems based on log-structured B-trees and on LSM-trees (Figure 7). Per-key updates are inserted into an in-memory index, or MemTable, and appended to a write-ahead log (WAL) for durability. The purpose of the MemTable, as in LSM-tree designs, is to buffer updates until we have reached a minimum batch size. Once this happens, the checkpoint pipeline sorts and de-duplicate key/value pairs in the MemTable, producing a batch which is applied to the most recent checkpoint (which is a TurtleTree). This process is repeated until $\chi$ batches (MemTables) have been applied, at which point the modified portions of the TurtleTreeare serialized and written to storage. Afterward, the WAL is trimmed.
Queries are processed by first consulting the MemTable, which represents recent updates (up to some size threshold). If the key is found here, the query finishes. Otherwise, the query proceeds to check recent MemTables in most-recent order, one per batch applied to the in-memory pending checkpoint. If the key is not found in any of these, the most recent durable checkpoint is searched.
Implementation & parameters We implemented TurtleKV in approximately 30k lines of C++ code. TurtleKV’s static parameters include $B$ (the block size), $M$ the total memory budget, the filter bit rate ($FBR$), and the number of threads involved in various stages of updating Memtables and the checkpoint pipeline ($T$). In this work, we investigate the value of dynamically changing checkpoint distance ($\chi$) on demand, This parameter varies the delay between creating pages or batches and writing them to disk.
TurtleKV needs to address two high-level challenges. First, it must guarantee high throughput and low latency. Second, for $\chi$ to have value, it must guarantee that I/O, not memory or CPU is the bottleneck. We next describe TurtleKV’s key components in more detail and how we address these challenges for them.
Figure 7. High-level Architecture of TurtleKV
4.1. TurtleKV Component Details
Write-Ahead Log. TurtleKV writes all key updates to a shared change log file initially, to provide data durability in the event of program termination before the next checkpoint update. This file is divided into fixed-sized blocks, which internally use a slotted page format to store variable-sized per-key edits.
To avoid low performance due to contention when logging updates, we implement a separation-of-concerns strategy where each user thread writes key updates to a thread-local block buffer to minimize contention. A single background thread is responsible for flushing data to storage. The background flusher thread uses a lock-free, periodic polling scheme to steal ownership of buffers from foreground threads. The polling period can be adjusted to trade off between minimizing the latency for updates to become durable, and the throughput achieved by update batching.
Change log blocks may be configured large enough (e.g., ¿ 4kib) so that writing to storage is not atomic. Therefore the block header includes a hash of its contents. At recovery time, any block whose hash does not match is assumed to be partially written, and is ignored. Each update to the change log is assigned a unique, monotonic sequence number using an atomic counter. This can be used during recovery to maintain a consistent order of events across restarts.
MemTables. We use Adaptive Radix Trees (ARTs) as the core data structure for indexing buffered updates (Leis et al., 2013). This allows TurtleKV to maintain an ordered index for range queries, providing low latency. A concurrent hash table index can also be enabled at runtime to further speed point queries, at the cost of slightly higher insertions since the ART is still used.
To mitigate contention between multiple writers, TurtleKV uses optimistic lock coupling (OLC) (Leis et al., 2016). We also exploit the fact that MemTables do not need to support true key removals, since deletions must be merged into the checkpoint tree. This allows us to use a simple arena-based memory allocation scheme, with each thread allocating memory from the system in 64kib chunks which are then divided as needed into ART nodes. TurtleKV MemTables rely on WAL block buffers to store actual key/value data, further simplifying the job of the ART index, since it does not need to copy update data.
Checkpoint Pipeline. We use the parallelism of the underlying hardware to perform the needed work. A foreground thread first places finalized Memtables in a queue, which is consumed by a background thread. This thread sorts and compacts key/value pairs to produce new batches of updates. Each batch is placed in another queue which is consumed by the thread responsible for updating the checkpoint tree.
The checkpoint pipeline must also manage copying of data as sorted key/value pairs are merged/compacted within buffer levels and flushed down the tree. By increasing $\chi$, we expect incoming key updates to be pushed further down the tree before they are written to storage. If we naïvely copy key/value data into page buffers during this process at every step, most of this work will be wasted in building pages which are never written.
To address this challenge, Checkpoint tree updates defer the copying of key/value data until the last batch has been applied (according to the $\chi$ parameter). Merge/compactions on in-memory checkpoint tree update buffer levels and leaves therefore operate on pointers, to avoid data copies. The individual merge/compact operations are parallelized using a background thread pool. This allows write-heavy workloads that execute on only a few threads to take advantage of available CPU cores when possible.
When the final batch for a checkpoint has been applied, all new pages (leaf, node, and filter) are built concurrently using the same thread pool used for merging. Once all pages have been built, the checkpoint is handed via a third queue to the final thread, which writes the page data in parallel to storage. Page lifetime is controlled via reference counting. The reference counts, as well as root references to subsequent checkpoint trees, are appended to a write-ahead log using a two-phase commit protocol to ensure crash consistency.
Once a new checkpoint has been made durable, the checkpoint field of the KVStore is updated and the finalized MemTables representing the batches used to create the new checkpoint are pruned from the Deltas stack. Pruning the MemTables from Deltas allows the change log to be trimmed, releasing storage resources and ensuring that recovery times do not grow unbounded.
Page Cache. TurtleKV uses LLFS (Astolfi et al., 2024), MathWorks’ Low-Level File System library. This means that storage resources are organized into different pools (extents) of storage, one per page size. [^7] Typically one for nodes (4kb), one for leaves/segments (2mb-16mb), and one for filters (16kb-256kb). Each of these pools is represented in LLFS as a PageDevice.
While it is advantageous to use large (>2>2mb) page sizes for leaves when writing new checkpoint pages, reading entire leaf pages while processing a point query can introduce a severe latency penalty. To address this issue, we add a kind of virtual PageDevice to LLFS, called a sharded view. Sharded views assign page identifiers to slices of larger pages; e.g., a 4kb sharded view of a 16mb page device allows the loading/caching of 4kb aligned slices. This feature allows the point query code path to load and cache only the parts of a leaf page that are needed to answer a query.
Using sharded views allows the TurtleKV’s large leaf page design to perform competitively in low cache configurations when compared to the alternate Bϵ+-tree design in SplinterDB, which uses 24mb B+-trees composed of 4kb pages instead of a single large page. TurtleKV’s sharded page view design has the additional advantage of being able to answer key presence queries (return true or false depending on whether the key is present, instead of returning its value) more efficiently, since key and value data are stored in separate parts of a leaf.
Caching policy. Even if we optimize all the other steps so that write I/O is the bottleneck, we still need to achieve a high cache hit-rate when updating the TurtleTree checkpoint. This is because the I/O improvements we obtain from increasing $\chi$ may be offset by increased I/O from reading pages containing data to merge, compact, or flush. To address this issue, we use use a modified CLOCK (Jiang et al., 2005) cache replacement policy where each page has a priority value (instead of a single bit) which is decremented when the clock hand passes it. This allows TurtleKV to provide information to the lower layer about which pages to make an extra effort to keep in memory.
5. Evaluation
We claim that TurtleKV demonstrates how a key-value storage engine can be built to support low cost re-tuning at runtime, with a dynamic trade-off range that provides excellent resource utilization and competitive performance for write-heavy, read-heavy and mixed workloads. The goal of the experiments in this section is to support this claim.
We aim to answer the following questions:
  • Question 1: Does TurtleKV offer competitive throughput, write amplification, and space amplification for write-heavy workloads?
  • Question 2: Does TurtleKV offer superior latency and throughput in read-heavy workloads, dominated by point and range queries?
  • Question 3: What are the properties, both in cost and benefit, of runtime adjustments to checkpoint distance to select an optimal tuning point in response to workload?
Methodology. To answer these questions, we use the Yahoo Cloud Serving Benchmark (YCSB), comparing TurtleKV to other state-of-the-art key-value engines, each representing a different optimization target: RocksDB, a write-optimized and widely used LSM-tree-based system; and SplinterDB, a state-of-the-art hybrid-optimized key-value store. Splinter and Rocks are run mostly in their default configurations; direct I/O was enabled in RockDB in order to control the size of the block cache (this resulting in a slight improvement in overall throughput).
In our benchmarking, we use two record sizes, 128 bytes and 1024 bytes, 8 bytes of which is the key. Each workload is preceded by an initial load phase, where we populate the database with 48GiB of data, 400M 128-byte records or 50M 1024-byte records. We test both out-of-cache and in-cache workloads; for out-of-cache, we limit cache size to 16GiB, $1/3$ of the pre-loaded data size, and for in-cache we allow up to 48GiB of cached pages. In our parameter selection process, we found that for out-of-cache workloads, a relatively small number of threads is needed to reach the point where I/O becomes the bottleneck. Correspondingly, for in-cache, the bottleneck shifts to CPU and memory bandwidth. Consequently, our out-of-cache tests use 4 client workload threads, and in-cache tests use 16 threads. We also include results for single-threaded tests to show latency of sequentially issued requests.
Hardware setup. Experiments were run on an AMD Ryzen Threadripper 7970x workstation (32 cores/64 threads), with 128 GB of DDR5-6000 registered ECC memory. The SSD used was an Intel P4800x 1.5TB, with 512 byte LBA configured, formatted with XFS. The P4800x is a data center grade Optane device supporting PCIe 3.0, rated for 2.4 GB/s read bandwidth and 2 GB/s write (550k/500k IOPs respectively). The system uses Ubuntu Linux 22.04 with kernel version 6.3. The CPU, while physically packaged as a single-socket part, was configured in the firmware (ASRock TRX50WS) to report its topology as a 4-node NUMA processor, one node per chiplet, each with 8 cores/16 threads.
(a) 400M $\times$ 128B, 1 thread
(b) 400M $\times$ 128B, 4 threads
(c) 400M $\times$ 128B, 16 threads
(d) 50M $\times$ 1024B, 1 thread
(e) 50M $\times$ 1024B, 4 threads
(f) 50M $\times$ 1024B, 16 threads
Figure 8. YCSB Results (higher is better): TurtleKV matches or exceeds performance of other engines on small records, and is competitive on large records. For 1024B records, TurtleKV performs better in higher cache configs due to separation of key and value data in large (16mib) leaf pages degrading query performance under cache pressure.
(a) 400M $\times$ 128B, Write Amp
(b) 50M $\times$ 1024B, Write Amp
(c) 400M $\times$ 128B, Space Amp
(d) 50M $\times$ 1024B, Space Amp
Figure 9. Write & Space Amplification (lower is better): TurtleKV has slightly higher WAF than SplinterDB, but uses less than half the space on disk.
(a) 400M $\times$ 128B, Workload E (scans)
(b) 50M $\times$ 1024B, Workload E (scans)
Figure 10. TurtleKV’s balanced compaction policy allows it to achieve high throughput when performing short range queries (<<100 records).
[…]
Figure 11. Checkpoint distance ($\chi$) tuning reduces writes and put latencies (a, e), increases upsert throughput, and increases memory usage (b, f). Adjusting $\chi$ at runtime is transparent (c, g) and cheap, costing only a few seconds to flush the current checkpoint (d, h).
5.1. Experiments
To evaluate write performance, we run YCSB Load (insertion of unsorted records into an empty database), Workload A (50%/50% mixture of reads and writes), and Workload F (50% reads and 50% read/modify/write). We measure and compare throughput (operations/sec), write amplification, space amplification, and memory usage. Generally, we are looking for high insert/update throughput (with correspondingly low average latency), low write and space amplification, and peak memory usage generally in line with the configured cache size settings.
To evaluate read performance, we run YCSB Workloads B (95% reads, 5% updates), C (100% point queries), and E (95% scans from 1 to 100 keys in length). We expect that the level-tiered structure of TurtleTrees will provide a performance boost in read-heavy workloads.
We explore the dynamic tuning capabilities of TurtleKV via a modification to the YCSB-C benchmark implementation. Before and after each workload, database drivers are given access to the workload properties and allowed to modify their dynamic configuration in response. This mechanism, while not realistic per se, allows an investigation into the possibilities offered by on-demand cache size and $\chi$ adjustment. We leave the problem of automatically detecting and adapting to workload characteristics using $\chi$ tuning for future research, for the moment focusing on the properties of the tuning knob itself.
To evaluate the effect of $\chi$-tuning, we run the YCSB Load with various values of $\chi$, from 1 to the memory limit of the test machine, followed by workloads C (100% gets) and E (95%/5% short scans/puts). The $\chi$ parameter is reset to 1 after the initial load and before the read workloads. This causes the checkpoint update pipeline ($\S$4.1) to be flushed, writing a new checkpoint to disk, trimming the WAL, and dropping old MemTables. We look to demonstrate that write amplification scales proportionally to put latency through the entire dynamic range, for large and very small (64 byte) records, and that the theoretically derived cost reduction of $O(-\log{\chi})$ is observed in practice. We also aim to show that the latency of re-tuning, i.e. of flushing the last checkpoint, is relatively minor and scales with $\chi$.
5.2. Results
Overall we found our experimental data supports the claim that TurtleKV offers performance matching or exceeding state-of-the-art key-value engines, with the flexibility of dynamic tunability. Below, we present the key results of our evaluation.
[…]
TurtleKV uses less than half of the on-disk space that SplinterDB does for the same data. Turtle’s write throughput trails Splinter’s in some cases because it has higher write amplification. How much checkpoint distance could reduce this was limited by the available memory. If the overhead of maintaining in-memory, unflushed checkpoints can be reduced, our approach might be able to surpass the highly write-optimized, lazily compacted design of Splinter. This potential warrants further research.
[…]
Result 3: Dynamic checkpoint distance based tuning offers dramatic improvements in write performance, but its memory cost yields diminishing returns in the limit. Our data supports the theoretical analysis of the benefits of $\chi$-based tuning. Tuning $\chi$ to the limit of available memory allows write amplification to be reduced from 12-14 down to 2-4, a dramatic improvement (Figure 11a, e). As claimed, $\chi$-based write optimization carries no residual penalty (Figure 11c, g), as shown by the stable query throughput curves overlaid by increasing insert throughput. The re-tuning cost (latency) of $\chi$, which is dominated by the flushing of a partial checkpoint, is under 10 seconds for most values, though it does increase with checkpoint distance (Figure 11d, h). The limitation of $\chi$ is that it does offer diminishing returns at higher memory usage levels (Figure 11b, f). This presents an opportunity for future research.
6. Related Work
The theoretical trade-off space constraining external memory indexing is characterized by Brodal and Fagerberg (Brodal and Fagerberg, 2003) for comparison-based dictionaries, and by Conway et al. (Conway et al., 2018) for hash-based dictionaries. The key insights of these works are echoed and elaborated in practical terms by the RUM Conjecture (Athanassoulis et al., 2016).
Recent efforts to explore trade-offs in LSM-tree based systems include Monkey (Dayan et al., 2017), a key-value store which optimizes queries through an analytically optimal allocation of memory to Bloom filters at various levels of the tree; Dostoevsky (Dayan and Idreos, 2018), which further improves space-time trade-offs for LSM-based systems by varying compaction policy from lazy at lower levels to greedy at higher ones; Wacky (Dayan and Idreos, 2019), which introduces adaptive level sizing and finer-grained control over compaction policy to even better optimize the trade-off curve; and Cosine (Chatterjee et al., 2021), which builds on the previous three and uses machine learning to optimize data structure design for specific cloud-based key/value storage workloads. We note that Dostoevsky and Wacky in particular share some thematic similarity to TurtleKV’s scheme of optimizing writes without harming future query performance by deferring on-disk compaction of recent updates and compacting older data more aggressively. Endure (Huynh et al., 2022) tunes size ratio, memory allocations, and compaction policy to maximize throughput under workload uncertainty.
Notable work in extending the trade-off tuning potential of B-tree based stores includes the Bw-Tree (Levandoski et al., 2013b; Levandoski and Sengupta, [n.d.]; Wang et al., 2018), a log-structured B-Tree which uses lock-free techniques to achieve low overheads for in-memory operations in high concurrency settings; Bf-Trees, a similar approach which introduces the innovation of caching mini-pages using a circular buffer to reduce I/O amplification while minimizing the overhead of cache maintenance; and the Bwe-Tree (Wang et al., 2024), which describes many practical improvements to the Bw-Tree design to optimize for fast SSDs. LSB-Trees (Kim and Lee, 2015) propose an alternate mechanism for storing B-Tree deltas using a combination of physical and logical logging. The relationship of Log-Structured B-Trees to the overall continuum of trade-offs in the external memory dictionary design space is explored by Idreos et al. (Idreos et al., 2019).
Several recent works address key/value engine design in the face of very fast modern storage devices. SplinterDB (Conway et al., 2020) and Mapped SplinterDB (Conway et al., 2023) introduce innovations in the design of external memory dictionaries, extending Bϵ-Trees to STBϵ-Trees by applying ideas from size-tiered LSM-trees to achieve high update efficiency without excessively penalizing point queries. FASTER (Chandramouli et al., 2018a, b) combines log-structured storage with in-memory concurrent hash tables to scale point-query-only based workloads to modern SSDs, but lacks first-class support for range operations. KVell (Lepers et al., 2019), later extended to KVell+ (Lepers et al., 2020), achieves high performance on modern SSDs using a shared-nothing design based on independent sharded indices maintained on separate threads. KVell only maintains key-based indices in memory (not in storage), trading scalability and recovery time for lower space-/write-amplification.
7. Conclusion & Future Work
With TurtleKV, we introduce a novel approach to read and write optimization that dynamically trades memory to establish a desired trade-off point on demand. The TurtleTree design and its accompanying tuning knob, checkpoint distance ($\chi$), achieves this property by creating an unbiased on-disk structure whose read and write I/O cost can be directly tuned via dynamic memory allocation without requiring external data structures that permanently consume available memory resources.
This dynamic approach to performance optimization opens many paths for future research and improvement of key-value storage techniques. These include extending the dynamic range of $\chi$-based tuning by further optimizing in-memory checkpoint tree updates, reducing memory requirements through the use of streaming compaction and other techniques, and exploring more of the Bϵ+-Tree design space to discover data structures beyond the TurtleTree.

### v2

@@
 1. Introduction
-Key-value storage is a foundational building block upon which many data-management applications are built.
-Though key-value storage’s dictionary-based interface flexibly supports a diverse range of applications, concrete implementations must trade off between the cost of reads, writes, and memory usage.
-Prior work hypothesizes that only two of these dimensions can be optimized at the cost of the third (Brodal and Fagerberg, 2003; Athanassoulis et al., 2016).
-The result is a proliferation of implementations tuned for specific workload characteristics (Meta Platforms, Inc., 2022; Conway et al., 2020; Conway et al., 2023; Apache, 2025; Google, [n.d.]).
-This growing diversity complicates database management: organizations must often deploy multiple databases to support different workloads and rely on expensive Extract-Transform-Load (ETL) pipelines to move data (Databricks, [n.d.]).
-Figure 1.
-Today’s key-value engines (left) trade reads directly against writes, then compensate for lost performance by spending memory.
-In contrast, TurtleKV(right) uses a balanced on-disk data structure.
-It dynamically trades off memory for increased write or read performance.
-With respect to the tradeoffs, state-of-the art databases limit memory usage and choose between reads or writes for the remaining dimension they wish to optimize (left diagram in Figure 1).
-They use on-disk structures biased for performance in this remaining dimension and use other mechanisms to recover lost performance in the unoptimized dimension.
-For example, SplinterDB (Conway et al., 2020; Conway et al., 2023) uses a on-disk data structure, STBϵ-trees, biased toward writes and uses a limited amount of memory for quotient filters/maplets to improve read performance.
-TreeLine (Yu et al., 2022) uses update-in-place grouping mechanisms and spends memory on histograms to improve reads.
-Industry approaches, such as RocksDB (Meta Platforms, Inc., 2022), optimize for writes via LSM-trees and use Bloom filters to improve read performance.
-These approaches are static.
-They find a single read/write optimization point and permanently change the on-disk data structure, creating a residual memory cost for the entire lifetime of the data.
-Though powerful, the above approaches curtail the maximum achievable performance in the unoptimized dimension.
-As a consequence, supporting load-then-read workloads with high performance requires either spending significant I/O changing the on-disk format or using an expensive ETL pipeline to move data between different databases.
-We present TurtleKV, a key-value engine that uses an unbiased on-disk structure suited for reads and writes.
-It includes a tuning knob that dynamically controls the allocation of memory resources on demand within a fixed budget to improve read or write performance as needed (right diagram in Figure 1).
-The knob allows memory usage to be re-allocated to the alternate dimension (e.g., reads) when high performance in currently-chosen dimension is no longer needed (e.g., writes).
-TurtleKV is made possible by the following two observations.
-First, though the trade-off between memory, reads and writes may be fundamental, the choice of which dimensions to trade-off against are not.
-Second, many common workloads do not need high write and read performance simultaneously, so memory can be fungibly exchanged between these two dimensions as needed.
-For workloads that require high read-and-write performance simultaneously, TurtleKV allows human operators to adjust the tuning knob in a weighted way to reflect their operational goals.
-To support TurtleKV, we introduce a new external memory dictionary data structure, the TurtleTree, which improves read performance and bounds worst case write I/O behavior.
-TurtleTrees are a fusion of B+-trees and LSM-trees, leveraging the strengths of both to achieve an unbiased state, balanced between read and write performance.
-B+-trees, which offer high read performance, provide the high-level structure with data organized in tree-node pages.
-LSM-trees, which offer higher write performance, are used within the tree-nodes to buffer updates.
-By carefully managing the TurtleTree structure, we can control the rate at which updates flow down the tree, bounding the number of I/Os required for reads and writes.
-This property results in a balanced on-disk structure that can be dynamically tuned for greater read or write performance.
-We identify TurtleTrees as belonging to a previously unidentified category of data structures, which we name Bϵ+-trees.
-Such Bϵ+-trees differ from Bϵ-trees and other buffered B-trees by adding a level of indirection between tree-node pages and pages containing per-node buffered updates.
-SplinterDB’s size-tiered Bϵ-trees (STBϵ-trees) also belong to the Bϵ+-tree family.
-TurtleTrees differ from STBϵ-trees in that they use level-tiered LSM-trees instead of size-tiered ones.
-This change shifts the focus from write performance to more unbiased point between reads and writes, while also allowing for more fine-grained control over write amplification.
-In contrast to traditional level-tiered LSM-trees, TurtleTree’s design limits write I/O costs by bounding the number of pages flushed to disk at any time.
-Furthermore, TurtleTrees and other data structures in the Bϵ+-tree family are complementary to existing read/memory trade-off mechanisms such as page caching and AMQ filtering because of the separation of data (leaf pages and buffer segments) from metadata (interior nodes and any filters).
-Checkpoint distance ($\chi$), which is TurtleKV’s dynamic tuning knob, controls the portion of TurtleKV’s total memory devoted to improving write performance and can achieve the performance of write-optimized, size-tiered approaches on demand.
-Concretely, this knob sets the number of in-memory batch updates that can be performed on a TurtleTree before updates are written to disk and combines synergistically with TurtleTree ’s control over flushed pages.
-A higher value allows more data to be merged in memory, reducing write amplification and thus increasing write performance.
-To dynamically improve read performance when under cache pressure, TurtleKV can choose a smaller value, so that memory that is unused for writes maybe (re-)allocated to cache more data pages as needed.
-The key benefit of this knob is that unlike existing systems that permanently spend memory by creating new structures such as filters, tuning decisions are not anchored to on-disk structures at the point data is written.
-TurtleKV can dynamically optimize reads or writes by simply adjusting the allocation of memory.
-[…]
-Key contributions:
-  1. 1)
-    An exploration of the benefit of trading off memory dynamically for increased write or read performance.
-The design of an architecture that enables high dynamic tuning range by starting with a unbiased on-disk data structure and dynamically trading memory for performance.
-  2. 2)
-    TurtleTrees, a novel on-disk data structure that merges the benefits of read-optimized B+-trees with write-optimized Level-tiered LSM-trees.
-The identification and formal definition of an important class of data structures to which TurtleTrees belong, Bϵ+-trees
-  3. 3)
-    The concept of checkpoint distance ($\chi$), which allows more batched updates to reside in memory before being written to disk.
-This limits write amplification as previous batches can expire in memory before being written.
-  4. 4)
-    TurtleKV, a key/value storage engine that uses TurtleTrees and checkpoint distance to achieve high levels of performance on modern SSDs.
-2. Overview
-Brodal and Fagerberg’s Lower Bounds (Brodal and Fagerberg, 2003) and the RUM Conjecture (Athanassoulis et al., 2016) show that key/value stores are constrained by a three-way trade-off between main memory usage, read operations, and writes.
-This means that storage engines have at most two degrees of freedom when selecting a tuning point for a given application: for example, committing to a cost function for reads and memory forces a minimum cost for writes.
-A key insight of our work is that, while this trade-off is fundamental, the mechanisms in common use by today’s storage engines to navigate the RUM space are not.
-This means it is theoretically possible to devise a dictionary type affording mechanisms which trade, for example, writes against memory independent of reads, and reads against memory independent of writes.
-The goal of our work is to demonstrate mechanisms which do this in a way that avoids coding a given tuning point into the structure of stored data.
-This allows tuning points to be chosen dynamically, by re-allocating memory in different ways to optimize for the workload at hand, expanding the range of applications which can be supported by a single general-purpose key/value store.
-In this section we give a brief review of current practice with respect to trade-off tuning in general purpose key/value storage engines, grouped by tuning goal: read-optimized, write-optimized, and hybrid..
-We then present the key ideas of this work: Bϵ+-trees, TurtleTrees, Write Optimization via Checkpoint Distance, and On-Demand Optimization.
-2.1. Current Approaches to Trade-off Tuning
-Here we describe and critique the major strategies presently in use to navigate the RUM trade-off space based on data structure type.
-Figure 2. Bϵ-tree $\rightarrow$ Bϵ+-tree structural transformation
-Read-Optimized (B-trees).
-B-trees (Comer, 1979) provide optimal (Yi, 2012) read performance and can utilize memory to reduce the number of reads required to process queries by caching pages containing tree-nodes.
+Today the builders of complex, cloud-based applications and systems have many different key-value stores to choose from, each with their own biases and trade-offs (e.g., read-optimized vs write-optimized; online-vs-analytic, in- vs out-of-memory, ordered-vs-hashed, domain-specific-vs-general-purpose, etc.)(jermaineNovelIndexSupporting1999; chaBlinkHashAdaptive2023; giladEvenDBOptimizingKeyvalue2020; pelkonenGorillaFastScalable2015; ashkianiGPULSMDynamic2018; wangLavaStoreByteDancesPurposeBuilt2024; sollezaMachPluggableMetrics2022; renSlimDBSpaceefficientKeyvalue2017a).
+These are frequently difficult to tune to accommodate evolving requirements after a system is built (RocksDBTuningGuide; caoCharacterizingModelingBenchmarking; SmallDatumImpact2022; jagadeesanWorstCaseAverageCaseAnalysis2020; RevisitingB+treeVs2022; nguyenCharacteristicsMongoDBTrimbased2016; andreoliInducingHugeTail2023; fedorovaPerformanceComprehensionWiredTiger2018).
+As a consequence, we see a proliferation of multi-stores with complex and expensive logic for synchronization, data transfer, and invariant maintenance (etl_pipeline; kumarValuePropositionETL; mukherjeeComparativeReviewData2017; el-sappaghProposedModelData2011; limSILTMemoryefficientHighperformance).
+Engineering effort must be spent multiple times for development, debugging, optimization, and maintenance of each storage engine.
+One would not be without reason believing that a multitude of storage engines is inevitable.
+The RUM Conjecture (athanassoulisDesigningAccessMethods2016) posits a fundamental, three-way trade-off in the design of data storage/access methods between the resource overhead of reads (R), updates/writes (U), and memory/storage (M).
+If correct, this might seem to imply that the divergence of requirements for different workloads and applications means key-value stores inhabit a problem space where one size does not, and cannot, fit all(idreosKeyValueStorageEngines2020).
+Our work assumes the correctness of the RUM Conjecture as its starting point, but seeks to challenge this implication and propose an alternative approach to key-value storage engine design in the face of inherent trade-offs.
+The notion of a three-way trade-off in the RUM Conjecture means that at least two trade-off selection mechanisms, or tuning knobs, must be utilized to fully cover the space of optimal possibilities; e.g.
+reads-vs-writes (RW), reads-vs-memory (RM), or writes-vs-memory (WM).
+Our key insight is that these basic trade-offs have important differences, and that these differences make some inherently better suited for exploitation as tuning knobs.
+For example, if a suboptimal tuning point is selected for the read-memory trade-off, it is likely easy to correct because of the mechanisms involved: simply modify the cache size to decrease either the I/O cost of queries or the memory footprint of the system at the expense of the other.
+In contrast, if a system is improperly tuned with respect to the read-write trade-off, things aren’t so simple.
+This is because read-write tuning is largely accomplished via selection of primary data structure, which affects implementation (code) and the invariants of stored data, both of which are expensive to modify post-facto.
+While some dictionary data structures do support parameter-based read-write trade-off tuning (e.g., compaction policy in LSM-trees with adjustable fanout (oneilLogstructuredMergetreeLSMtree1996; luoLSMbasedStorageTechniques2020) ), when analyzed closely and evaluated empirically, we find the mechanisms in common use to be sub-optimal.
+In any case, even if the primary data structure of a system does support smooth, optimal tuning between read and write optimization (e.g, Bϵ-trees’ $\epsilon$ parameter (brodalLowerBoundsExternal2003; benderIntroductionBetreesWriteOptimization2015)), bias is still encoded at the storage level, making re-tuning expensive.
+Since write-optimization is very important to certain workloads (caoCharacterizingModelingBenchmarking), write-memory (WM) tuning must be supported as a first-class feature if we wish to avoid the downsides of RW tuning bias in stored data structures.
+To do this it is clearly necessary, though insufficient, to simply expose a tuning knob.
+The choice of data structure and the overall design of the system must work together to maximize the effectiveness of this trade-off.
+A good example of how data structures can anticipate a resource trade-off (RM) in a system design context is the pervasive use of B+-trees in preference to B-trees (comerUbiquitousBTree1979; bayerOrganizationMaintenanceLarge1970).
+By storing all key-value records at the bottom level of the tree, B+-trees offer increased branching factor and the possibility for all interior nodes to be cached in memory, reducing read amplification to $O(1)$.
+A key idea of the work presented here is to apply a similar principle in data structure selection to enhance the effectiveness of the write-memory (WM) trade-off, elevating it to first-class status in key-value system design.
+We present TurtleKV, a proof-of-concept for a new approach to key-value storage oriented around RM and WM tuning knobs which dynamically navigate the RUM trade-off space without having to change the structure of stored data.
+TurtleKV is built around a novel data structure, the TurtleTree, which offers two important properties: 1.
+it achieves a neutral, balanced compromise between the extremes of read- and write-optimized external memory dictionaries, and 2.
+it affords a smooth trade-off between memory usage and write amplification which is simultaneously efficient, scale-independent (in terms of data size $N$), conceptually simple, and wide-ranged in effect.
+TurtleKV implements a system architecture which is similar at a high level to conventional LSM-based key-value stores, with the addition of several crucial optimizations to eliminate in-memory bottlenecks so that reductions in write-amplification translate to increased throughput.
+Because TurtleKV starts from a fundamentally stronger premise, it avoids many of the complex workarounds and features which tend to creep into modern key-value stores in order to try to mitigate their intrinsic limitations.
+As a result, we show that TurtleKV outperforms today’s key-value engines, while offering superior tunability, with a modest code base of similar or lesser complexity.
+In this work we demonstrate that write-vs-memory tuning mechanisms are often not given first-class status in the design of current key-value stores ($\S$2.2).
+For example, while RocksDB (RocksDBPersistentKeyvalue), one of the most popular key-value storage engines today, supports runtime adjustment of write buffer size (RocksDBTuningGuide), and our experiments show this parameter is highly effective at spending memory to reduce write amplification, doing so yields very little benefit in terms of increased end-to-end performance.
+WiredTiger (WiredTigerWiredTigerDeveloper), another popular storage engine on the opposite end of the read-write spectrum, also supports WM tuning, and even demonstrates end-to-end benefit in terms of increased insert throughput over wide range of tuning points.
+However, WiredTiger’s B+-tree based design offers little benefit at low-to-moderate memory allocations due to the inherent limitations of the data structure (idreosDesignContinuumsPath2019).
+Finally, we find that SplinterDB (conwaySplinterDBClosingBandwidth2020; conwaySplinterDBMapletsImproving2023; SplinterDBb) which is optimized for high concurrency and fast storage devices, is able to take advantage of higher memory allocations to reduce write amplification and average insertion latency for randomized insertion workloads.
+However, since SplinterDB does not offer any way to express a preference for optimization target (i.e., use cache memory to optimize writes vs reads), in mixed workloads the optimization effect of higher cache size may become unpredictable, as the dualing concerns of updates and queries vie for limited memory resources.
+The contributions of this work are:
+  1. (1)
+    An analysis of how current key-value stores and their data structures fall short of fully exploiting the potential of the RUM space, especially with respect to the write-memory trade-off
+  2. (2)
+    A novel external memory dictionary type, the TurtleTree, within a previously unidentified class of data structures, Bϵ+-trees, which combines ideas from both read- and write-optimized dictionary types
+  3. (3)
+    A key-value storage engine design, TurtleKV, which uses TurtleTrees to offer simple and smooth trade-off tuning for a wide range of resource allocations
+  4. (4)
+    An experimental evaluation of TurtleKV as compared to SOTA key-value stores WiredTiger, RocksDB, and SplinterDB
+2. Read, Update, Memory: Pick Two
+Figure 1. RUM as a three-dimensional resource space
+The RUM Conjecture (athanassoulisDesigningAccessMethods2016) states that if one holds any of the three resource overheads (read, update/write, and memory/space) constant, the remaining two are placed in a zero-sum, trade-off relationship.
+Geometrically, this can be visualized, as in Figure 1, by a three dimensional resource space, where points represent the actual resource cost of a concrete data access method with respect to a particular workload.
+The notion of a three-way trade-off, in this conceptualization, is represented by a two dimensional surface or manifold embedded within the larger-dimensional space (dark gray).
+Fixed resource commitments can be thought of as two dimensional slices of this surface orthogonal to the axis of the fixed resource, with the resulting intersection defining the trade-off curve between the remaining two resources.
+Navigating the optional two-dimensional subset of the RUM space requires the use of exactly two mechanisms, or tuning knobs, which serve as a coordinate system within the sub-space.
+In this section we analyze our problem space by considering three tuning knob candidates: the read-write trade-off (RW), the read-memory trade-off (RM), and the write-memory trade-off (WM).
+These are explored through the perspectives of data structure ($\S$2.1) and system design ($\S$2.2).
+We conclude with a sketch of the key ideas behind TurtleKV ($\S$2.3).
+2.1. Data Structures: Trade-off Tuning
+The most consequential design decision for a key-value storage engine is its choice of primary data structure.
+Here we review the major categories of external memory dictionary, grouped by their RW trade-off optimization target (read-optimized, write-optimized, and hybrid/balanced).
+We highlight the essential insights of each data structure, how each approaches trade-off tuning, and their inherent limitations.
+2.1.1. Read-Optimized (B-trees).
+B-trees(comerUbiquitousBTree1979) provide optimal (yiDynamicIndexabilityOptimality2012) read performance and can utilize memory to reduce the number of reads required to process queries by caching pages containing tree-nodes (RM trade-off).
 B+-trees, the most commonly implemented B-tree variant, exploit this mechanism by placing key/value data in the leaves, allowing all but the bottom layer of the tree to be cached in main memory so that queries can be answered with a single read.
-State-of-the-art approaches to reducing the write cost of updates in B-trees employ the use of a write-ahead log (WAL) to ensure durability of updates, plus a scheme for applying updates to the tree in memory.
+State-of-the-art approaches to reducing the write cost of updates in B-trees employ the use of a write-ahead log (WAL) to ensure durability of updates, plus a scheme for applying updates to the tree in memory (WM trade-off).
 We refer to these as Log Structured B-trees.
 Their goal is to defer rewriting pages containing updated keys, in hopes of collecting enough updates per page to amortize I/O.
-Bw-trees (Levandoski et al., 2013a) do this through the use of delta records applied in-memory to the main tree using lock-/latch-free techniques.
-The closely related Bf-tree (Hao and Chandramouli, 2024) use an alternate strategy of caching and updating sub-page slices (mini-pages) instead full pages, which has additional benefits in terms of lowering the memory cost of caching ”hot” records.
+Bw-trees (bwtrees) do this through the use of delta records applied in-memory to the main tree using lock-/latch-free techniques.
+The closely related Bf-tree (bftree) uses an alternate strategy of caching and updating sub-page slices (mini-pages) instead full pages, which has additional benefits in terms of lowering the memory cost of caching ”hot” records.
 Bw-trees and Bf-trees achieve good write efficiency either when memory is plentiful or when updates are clustered in a few sub-ranges of the key space.
-In the general case, however, Log Structured B-trees force a sub-optimal trade-off between memory and write amplification, especially as the total data size $N$ grows with respect to available memory $M$ (Idreos et al., 2019).
-Write-Optimized (LSM-trees).
-Log-Structured Merge (LSM)-trees (O’Neil et al., 1996) are a popular write-optimized alternative to B-trees.
-By deferring the merge of new updates with existing data, they maximize the amount of effective work achieved by each write operation.
-Unfortunately, this technique forces read costs (and space requirements) to increase, since it introduces the possibility that updates for a given key may be stored in multiple locations.
-Page caching can be used to trade memory against reads, but this mechanism is less effective for LSM-trees than B-trees because there is potentially much more data to cache.
-In addition to caching, two primary tuning mechanisms are employed by LSM-Tree systems: compaction policy and approximate memory query (AMQ) filtering.
+In the general case, however, Log Structured B-trees force a sub-optimal trade-off between memory and write amplification, especially as the total data size $N$ grows with respect to available memory $M$ (idreosDesignContinuumsPath2019).
+2.1.2. Write-Optimized (LSM-trees)
+Log-Structured Merge (LSM) trees (oneilLogstructuredMergetreeLSMtree1996) are a popular write-optimized alternative to B-trees.
+By deferring the merge (or compaction) of new updates with existing data, they maximize the amount of effective work achieved by each write operation.
+This results in a dramatic reduction in write amplification, but presents two challenges: first, it increases read costs (and storage footprint) by introducing the possibility that updates for a given key may be stored in multiple locations.
+Second, since the increasingly high bandwidth of modern storage devices favors performing merges in large sequential runs to maximize I/O utilization, update costs in LSM-trees tend to be amortized costs; i.e.
+tail latencies can be quite high.
+This motivates the need for effective de-amortized, or incremental, approaches to compaction.
+Page caching is less effective as an RM tuning mechanism for LSM-trees than B-trees because there is potentially much more data to cache.
+In addition to caching, two primary tuning mechanisms are employed by LSM-tree systems: compaction policy and approximate membership query (AMQ) filtering.
 Compaction policies divide into greedy (level-tiering) and lazy (size-tiering).
@@
 In a size-tiered LSM-tree, a tunable number $T$ of key duplicates are allowed to co-exist at a given level before their eventual merge into a larger level.
-Both compaction policies directly trade read cost against writes by altering the invariants of stored data.
+Compaction policy thus alters the invariants of stored data to achieve a RW trade-off mechanism.
 Approximate member query (AMQ) filters, e.g.
-Bloom Filters (Bloom, 1970), allow memory to be spent to reduce the read cost of point queries by having a dedicated filter (or part of a shared filter (Conway et al., 2023)) devoted to each location where a key might reside.
-AMQ filters offer a tuning parameter $\alpha$, the false-positive rate, which trades size against accuracy.
+Bloom Filters (bloomSpaceTimeTradeoffs1970), are an RM trade-off mechanism that reduce the cost of gets by having a dedicated filter (or part of a shared filter (conwaySplinterDBMapletsImproving2023)) devoted to each location where a key might reside.
+AMQ filters offer a tuning parameter $\alpha$ [^1] Also frequently named as $\epsilon$, the false-positive rate, which trades size against accuracy.
 [^2] A filter must always answer accurately if a query key is not in its set, but it may give a false positive result at rate $\alpha$.
@@
 This rules out, with probability $\prod_{i\in S}{\left(1-\alpha_{i}\right)}$, locations where the query will fail, preventing excess reads.
-If filters are used to maintain a target read cost in a size-tiered LSM-tree (with $T$ tiers per level), the total size of the filters must scale in proportion to $T$, since the total number of false-positives for a query is the sum of false-positives for all filters that pertain to a given key (Dayan et al., 2018).
-Therefore the $T$ parameter in a size-tiered LSM, together with $\alpha$, can be thought of as representing a trade-off between memory and write amplification.
-The problem with this mechanism is that the future read amplification one trades to achieve better write performance is reflected in the on-disk structure, and thus the memory cost for filters to compensate must be paid during the entire readable lifetime of write-optimized updates.
-Note that this is true even if the filter type (e.g.
-Quotient Maplets (Conway et al., 2023)) supports cache-friendly paging of partial filters into main memory, since the filter bit rate must nevertheless be increased to maintain a target query cost, regardless of how much of a filter is cached.
-In essence, write-optimization in the past generates a read/memory trade-off penalty in the future in such designs.
-We refer to this effect as the residual write-optimization penalty.
-Hybrid-Optimized (Bϵ-trees).
-Bϵ-trees (Bender et al., 2015; Brodal and Fagerberg, 2003) are a tunable external memory dictionary, able to trade reads against writes.
+If filters are used to maintain a target read cost in a size-tiered LSM-tree (with $T$ tiers per level), the total size of the filters must scale in proportion to $T$, since the total number of false-positives for a query is the sum of false-positives for all filters that pertain to a given key (dayanOptimalBloomFilters2018a).
+As $T$ increases, read amplification increases while write amplification goes down; as filter bits-per-key increases, memory usage rises while read amplification falls due to lower false-positive rate $\alpha$.
+It is theoretically possible therefore to coordinate the motion of these two parameters ($T,\alpha$) to target a fixed level of read amplification, while writes are effectively traded against memory (WM) using filter size-vs-accuracy as an intermediary.
+We refer to this strategy as filter arbitrage.
+2.1.3. Hybrid-Optimized (Bϵ-trees)
+Bϵ-trees (benderIntroductionBetreesWriteOptimization2015; brodalLowerBoundsExternal2003) are a tunable external memory dictionary, able to trade reads against writes.
 The basic B-tree structure is used as the foundation for a Bϵ-tree.
@@
 Similarly, decreasing $\epsilon$ to decrease the branching factor and increase the buffer size increases the height of the tree, which raises the cost of reading/caching a node-path to any given key.
-Size-Tiered Bϵ-trees, used by SplinterDB (Conway et al., 2020; Conway et al., 2023), are a recent re-imagining of the Bϵ-tree designed to close the ”bandwidth gap” between the capabilities of fast NVMe SSDs and traditional key/value storage engine designs.
-The main structural modification in STBϵ-trees is to add level of indirection between node pages and buffered data (Figure 2, 4 ).
-This change is analogous to the transformation between B-trees and B+-trees, and gives SplinterDB similar advantages in terms of being able to cache interior nodes for reading.
-SplinterDB adopts the strategy of selecting a single tuning point which attempts to offer excellent read and write efficiency by structuring per-node buffer data pages as a single-level, size-tiered LSM-tree-like structure, where each sorted run (or branch) is stored as a statically built B+-tree.
-Quotient Maplets, a variant of Quotient Filters, are used in this design to route point queries to the branches/tiers which may contain a given key.
-While benchmark performance data (YCSB) shows this to be a good compromise for point-query workloads on modern hardware, it nonetheless represents a trade-off between false positives leading to increased read amplification on the one hand, and the memory cost of maintaining maplets with sufficient bits per key to compensate for the number of possible redundant copies of a key stored at multiple tiers.
-The downside of this approach is felt most acutely in short (¡100 key) range scans, which must merge/compact a potentially large number of tiers to determine sequential keys.
+Figure 2. Bϵ-tree $\rightarrow$ Bϵ+-tree structural transformation
+Size-Tiered Bϵ-trees (STBϵ-trees), used by SplinterDB (conwaySplinterDBClosingBandwidth2020; conwaySplinterDBMapletsImproving2023), are a recent re-imagining of the Bϵ-tree designed to close the ”bandwidth gap” between the capabilities of fast NVMe SSDs and traditional key/value storage engine designs.
+The main structural modification in STBϵ-trees is to add level of indirection between the tree of interior node pages, or trunk, and buffered data, or branches (Figure 2, 4 ).
+This separation allows buffered updates to be flushed in larger batches than a traditional Bϵ-tree, enabling high I/O bandwidth utilization.
+It also enables flushes to be decoupled from compactions since only the branch references need to be changed, enabling SplinterDB’s novel flush-then-compact policy.
+Flush-then-compact has the added benefit of allowing updates with skewed key distributions to be compacted at a lower level of the tree, reducing write amplification for ”hot” records.
+However, due to their lazy compaction policy, STBϵ-trees must do more work to perform scan queries, which hurts performance.
+2.2. System Design: Write-Memory Tuning
+We now shift our attention from data structures to the concerns of system design, to explore how well current key-value storage engines support write-vs-memory (WM) trade-off tuning.
+We present three case studies of systems spanning the read-write (RW) spectrum covered in $\S$2.1: WiredTiger, RocksDB, and SplinterDB.
+All are high-performance key-value storage engines optimized for concurrent workloads run on hardware with multiple processor cores and solid-state storage devices.
+RocksDB and WiredTiger in particular are heavily relied upon in real-world systems, and have been the beneficiaries of a large investment in engineering effort.
+The main differentiator in these systems is the choice of primary data structure: WiredTiger uses B+-Trees [^3] WiredTiger also supports LSM-based indexing, but it is not the default., RocksDB is based on LSM-Trees, and SplinterDB introduces the novel STBϵ-tree.
+An investigation into the limitations of how each of these systems support WM trade-off tuning elucidates some of the central issues and provides motivation for the design of our prototype, TurtleKV.
+Specifically, we present three findings: first, a system’s ability to navigate the WM trade-off is highly dependent on the choice of primary data structure; second, the ability to reduce write amplification does not always translate into increased real-world performance; and third, supporting I/O vs memory trade-offs in general does not imply the ability to target specific tuning points in terms of memory-based read- or write-optimization.
+We summarize the key results of these case studies in Table 1.
+WiredTiger RocksDB SplinterDB TurtleKV
+Data Structure supports
+efficient WM trade-off?
+System Design has no
+in-memory bottlenecks?
+First-class write-memory
+trade-off tuning knob?
+Table 1. The three key requirements for dynamic write-memory trade-off tuning in key-value storage engine design.
+2.2.1. WiredTiger
+WiredTiger is the default storage engine for MongoDB.
+It is designed to support high levels of concurrency for transactional key-value workloads.
+To accomplish this, it uses lock/latch-free algorithms and multi-version concurrency control (avoiding update-in-place) to implement its primary data structure, a B+-tree.
+Updates to the B+-tree in WiredTiger take place in ”dirty” buffers pinned to its in-memory page cache, where new versions of existing pages (or entirely new pages) are built.
+Dirty pages are flushed to durable storage in response to a checkpoint request or when their total size exceeds a runtime-configurable threshold444eviction_dirty_target for background threads, eviction_dirty_trigger for application/foreground threads..
+Write-ahead logging can be enabled for durability of record-level updates in between dirty page flush triggers.
+Thus, WiredTiger supports dynamic write/memory trade-off tuning through its dirty page eviction (write-back) caching mechanism.
+Figure 3a shows how WiredTiger responds to increased memory allocation for dirty page buffers, running a workload with uniformly distributed keys inserted in random order.
+Here we see that average insert latency tracks with write amplification through the entire dynamic range of the resource trade-off, which is good.
+Reductions in write amplification factor (WAF), however, are minimal for 16mb $\leq M\leq$ 500mb; at higher memory allocations, WAF reduction accelerates.
+This is because, for random (non-skewed) key distributions, the expected benefit of buffering in B+-Trees is related to the total dictionary size $N$.
+If each page can hold $B$ items, then the expected write amplification $W_{x}$ for record $x$ is $O(\max(1,\min(\frac{N}{M},B)))$.
+Note, $W_{x}\in\{1,B\}$, since each record must be written at least once ($1$), and in the worst case it is re-written every time each other record in the same page is written ($B$).
+2.2.2. RocksDB
+RocksDB is used at Meta as the primary storage engine of ZippyDB (a distributed key/value storage service developed by and used within Meta), UDB (a MySQL storage layer for social graph data) (caoCharacterizingModelingBenchmarking), and by countless other open source projects and commercial systems.
+RockDB’s Tuning Guide (RocksDBTuningGuide) presents a formidable set of user-selectable policies and tuning knobs which optimize for various use cases.
+We identify two such knobs which appear to have the greatest impact on the WM trade-off: the size $\overline{M_{w}}$ of an in-memory table (MemTable) 555rocksdb.options.write_buffer_size, and the number of MemTables $|M_{w}|$666rocksdb.options.min_write_buffer_number_to_merge, rocksdb.options.max_write_buffer_number collected in memory before being flushed to the first level of external storage ($L_{0}$).
+We find that varying $\overline{M_{w}}$ with fixed $|M_{w}|$ is superior to varying $|M_{w}|$ for fixed size $\overline{M_{w}}$, due to the degradation of query performance as the number of MemTables increases; both approaches offer the same benefit in write-optimization.
+RocksDB’s default compaction strategy (leveled) uses incremental merging, in which levels are segmented into fixed-size sorted string table (SSTable) files.
+When the total size of all segments within a level $L_{i}$ grows too large, an SSTable (of size $\overline{M_{w}}$) within $L_{i}$ is selected to be compacted with data in the next larger/older level $L_{i+1}$.
+The compaction merges the single segment from $L_{i}$ with all segments in $L_{i+1}$ whose key range overlaps, producing a new sequence of SSTable segments whose keys are ordered and unique.
+The total write amplification $W_{x}$ for a given record $x$ can be calculated as the number of times $F$ (for fanout) the record is merged into a segment at a given level, times the number of levels in the LSM [^7] The reason there are $\log_{F}({N/M})=\log_{F}{N}-\log_{F}{M}$ levels for data set size $N$ is that the smallest level’s size is always at least $M$; one can think of the first $\log_{F}{M}$ merges of a record as happening in main memory, inside a MemTable.: $W_{x}=F\cdot\log_{F}({N/\overline{M_{w}}})$.
+Increasing MemTable size $M$ therefore spends memory to reduce write amplification, independent of total size $N$: $\Delta W_{x}=-O(\log{\overline{M_{w}}})$.
+(a) WiredTiger
+(b) RocksDB
+(c) TurtleKV
+[…]
+(a) SplinterDB
+(b) TurtleKV
+[…]
+Figure 3b shows the result of repeatedly running a uniform random insertion workload against an empty RocksDB database, while increasing $\overline{M_{w}}$ from 16mb up to 4gb.
+When 16mb $\leq\overline{M_{w}}\leq$ 1gb, we see the predicted logarithmic reduction in write amplification (x-axis is log scale).
+However, for MemTable/SSTable sizes over 1gb, the write optimization effect flattens out, since the inserted/flushed data never escapes $L_{0}$ before the workload completes [^8] Key ranges of SSTable files in $L_{0}$ are allowed to overlap, so as not to stall MemTable flushes; at least 4 files must be present in $L_{0}$ to trigger a compaction to $L_{1}$..
+Unfortunately this reduction in write volume does not translate into lower average insert latencies beyond RocksDB’s default value of $\overline{M_{w}}$ =64mb.
+This underscores the necessity of optimizing main-memory costs (CPU, memory bandwidth, and internal synchronization bottlenecks) if one targets I/O reduction as a tuning mechanism.
+2.2.3. SplinterDB
+SplinterDB largely adopts the strategy of selecting a single tuning point which attempts to offer excellent read and write efficiency by structuring per-node buffer data pages as a single-level, size-tiered LSM-tree-like structure, where each sorted run (or branch) is stored as a statically built B+-tree.
+Quotient Maplets, a variant of Quotient Filters (pandeyVectorQuotientFilters2021a), are used in this design to route point queries to the branches/tiers which may contain a given key.
+SplinterDB’s STBϵ-trees have configurable branching factor, which is tied to the number of tiers per level $T$.
+Quotient filter bit rate may also be tuned, and thus STBϵ-trees theoretically support a degree of write-memory tuning via filter arbitrage.
+[…]
+2.2.4. TurtleKV
+For comparison, we show TurtleKV’s results for both N=100M write-buffer size scaling and N=500M cache size-scaling workloads in Figures 3c and 4b respectfully.
+These results show that TurtleKV’s primary data structure, the TurtleTree, efficiently reduces write cost at a rate of $O(\log(M))$, where $M$ is the memory spent on write-optimization, like level-tiered LSM-tree-based RocksDB.
+However, because of the system design ideas elaborated in the following sections, TurtleKV is able to turn WAF reduction into higher performance over a very wide tuning parameter range, similar to WiredTiger.
+When configured for write optimization, TurtleKV keeps pace with SOTA engine SplinterDB, but because TurtleKV’s data-parallel strategy for in-memory compactions is able to achieve even load balancing across processors, its scaling curve is much smoother than SplinterDB’s (Figure 4).
 2.3. Towards Dynamic Tunability: Key Ideas
-In this work, we propose a dynamic, unbiased approach that achieves desired read and write performance on demand while avoiding the costly re-tuning or data movement required in existing, statically-tuned systems.
-Key Idea 0: Bϵ+-trees.
-A baseline requirement for dynamic optimization is an unbiased data structure with sufficient flexibility to achieve both high write and read performance.
-To achieve this property, we identify and build upon what we consider to be the core insight behind SplinterDB’s STBϵ-trees: moving buffered updates in a Bϵ-tree out of the node pages and into pages referenced from nodes (Figure 2, 3 $\Rightarrow$ 4 ).
-Following the observation that this transformation is quite similar to the difference between B-trees and B+-trees (Figure 2, 1 $\Rightarrow$ 2 ), we coin the term Bϵ+-tree to describe this new class of dictionaries.
-Moving update data out of node pages improves flexibility by offering potential advantages for both read- and write-optimization in Bϵ+-trees.
-Read efficiency benefits from smaller nodes being more easily cached; likewise, queries benefit from reduced tree height as a result of a higher branching factor, since the portion of a node devoted to pivots can be larger without sacrificing buffer capacity.
-In terms of write efficiency, being able to buffer more data per node allows updates to be moved down the tree in larger batches, allowing more work to be achieved per write during flushes.
-The identification of Bϵ+-trees opens the door for research into a set of important design dimensions around how the data within a node buffer is structured.
-For example, instead of the STBϵ-tree’s flat structure of references to static B+-trees with size-tiered compaction policy, adopting a more greedy (level-tiered) compaction policy across levels within a node shifts away from a write-optimized focus towards a more unbiased one.
-A key contribution of this work is to begin exploring some of this space of possibilities to better understand their potential.
-Key Idea 1: TurtleTrees.
-With Bϵ+-trees as a foundation, we introduce a novel data structure that leverages an unbiased, level-tiered node buffer structure and flush policy to enable dynamic read and write optimization.
-We name this new data structure the TurtleTree, shown in Figure 3.
-TurtleTrees are an instantiation of Bϵ+-trees in which buffer c -referenced data segments d are stored as flat leaf-like pages organized into a small, fixed capacity level-tiered LSM-tree per node a .
-Unlike size-tiered Bϵ+-trees, which are biased towards improving write performance at the cost of reads or log structured B-trees, which are biased towards improving reads over writes, the level-tiered TurtleTree sits at a middle point between these two poles.
-Like LSM-trees and other Bϵ-tree variants, TurtleTrees permit updates for a given key to be stored in multiple places (though keys must be unique within a given buffer level).
-By utilizing a level-tiered compaction rather than size-tiered, the number of potential locations is reduced, accelerating reads.
-TurtleTrees can also still use the various memory/read trade-off mechanisms discussed earlier, such as page caching and AMQ filtering, to further reduce read cost.
-In fact, we will see in this work that TurtleTrees can achieve read performance that is on par with or exceeds current state of the art read-optimized key/value stores.
+Table 1 expresses the three high-level goals which drive the design of TurtleKV.
+In this section we provide a brief sketch of the key ideas which address each goal.
+These ideas will be elaborated in the following sections.
+(a) Logical Structure
+(b) Page Layout
 Figure 5. The TurtleTree data structure.
-Setting aside reads, the question remains as to how to restore the write efficiency we lost by changing the intra-node compaction policy to level-tiered.
-We present a mechanism which achieves this in the next key idea (2).
-In preparation, we highlight a unique characteristic of TurtleTrees.
-By building a level-tiered structure within a Bϵ+-tree, we can bound the I/O cost of writing batches of modified data pages from the TurtleTree in-memory to disk, avoiding the write amplification penalty of traditional level-tiered LSM-trees.
-In fact, by setting the number of data pages d referenced by the buffer c roughly equal to the number of pivots b  [^2] The diagram in Figure 3 is conceptual; the exact ratios of pivots to buffer segments, and their relative sizes, are not accurate/to-scale., we guarantee that for each batch of pages inserted in the TurtleTree, we can flush an equal-sized batch to a lower level, bounding the growth of write amplification.
-This fine-grained control over how writes propagate down a TurtleTree allows us to choose exactly how many pages to flush and when to flush them.
-Key Idea 2: Write Optimization via Checkpoint Distance.
-Given the unbiased TurtleTree structure, we can now introduce checkpoint distance, a tuning knob that improves write performance in exchange for memory without affecting future performance via residual optimization penalties.
-When receiving a batch of updates, TurtleKV applies the batch to a TurtleTree in-memory and can immediately write the modified pages to disk, creating a checkpoint.
-However, it is also possible to defer this write, buffering the changes from multiple batches within the TurtleTree in-memory before creating the corresponding checkpoint on disk.
-We refer to the number of batches $\chi$ buffered in-memory as the checkpoint distance.
-By modifying $\chi$ at runtime, it is possible to trade write performance for memory in TurtleKV; specifically, the number of writes required per key update is reduced by $O\left(\log{\chi}\right)$!
-This reduction is independent of the total size $N$ of stored data, making it a viable tuning knob regardless of the on-disk size or structure.
-To understand why this is the case, consider the lifetime of one of the buffer data pages d shown in Figure 3.
-As new batches are applied to a node, the merging/compaction process causes data pages at higher levels to be frequently replaced with merged data and written to lower levels.
-This means that the closer a buffer segment’s level is to the root, the shorter its effective lifetime.
-By buffering batches in the TurtleTree, these short-lived data pages naturally expire and no longer need to be written to disk.
-Leveraging this property, we arrive at a key insight: by spending memory to defer writing these short-lived pages, we can avoid unnecessary writes and reduce write amplification without changing the stored structure of the eventual checkpoint.
-Prior work such as SplinterDB (Conway et al., 2020; Conway et al., 2023) or TreeLine (Yu et al., 2022) permanently spend memory by storing additional filters or histograms to index into their data pages, incurring a residual optimization penalty each time they want to optimize performance.
-In contrast, checkpoint distance leverages the TurtleTree structure to improve write performance with no additional memory cost beyond the tree itself and its internal metadata.
-This means that checkpoint distance can be adjusted up or down to control write amplification without incurring any residual optimization penalty.
-Key Idea 3: On-Demand Optimization.
-Instead of one-time, static optimization, TurtleKV leverages the TurtleTree structure to optimize both reads and writes on-demand, creating a general-purpose KV engine that negates the need for expensive data movement present in current Extract-Transform-Load (ETL) pipelines.
-TurtleTrees offer a tunable external memory dictionary which can navigate the RUM trilemma space using only read vs memory and write vs memory tuning mechanisms.
-We take advantage of this property by using checkpoint distance as a tuning knob to dynamically determine how much of the total available memory to spend on optimizing writes or reads.
-When the checkpoint distance is larger, more memory is spent buffering batches within TurtleTrees, optimizing writes, while when the checkpoint distance is smaller, more memory is dedicated in the page cache for data pages, optimizing reads.
-Since the tuning mechanism trades reads and writes against memory rather than reads vs writes directly, adjusting the memory allocation does not affect the underlying data structure.
-Thus, TurtleKV can dynamically adjust the memory allocation without permanently affecting either read or write performance, allowing for on-demand optimization of reads or writes depending on application requirements.
-In systems whose tuning mechanisms trade reads against writes directly, such as compaction policy in LSM-tree storage engines, the cost of re-tuning is high because it involves changing the stored structure of data to target a new trade-off point.
-Historically, applications have implemented this kind of ”tuning by moving” strategy through the use of multiple data stores and Extract-Transform-Load (ETL) pipelines, which transfer data between storage systems to optimize it for various application requirements.
-The fact that ETL and multi-stores have been quite popular despite high cost and increased system complexity speaks to the need for data management systems which can re-tune on demand to support changing workloads.
-In contrast to expensive data movement, TurtleKV can simply spend memory whenever it is advantageous to do so to accelerate read- or write-heavy workloads: pay now, not later.
-3. TurtleTrees & Checkpoint Distance
-[…]
-Compatibility with other mechanisms.
-We focus on $\chi$ as the main parameter to improve TurtleTrees’ performance.
-TurtleTrees are also compatible with other mechanisms to increase read performance: increasing page-cache size and read-filter bitrate.
-In this work, we assume static sizes for these additional parameters to isolate the effect of TurtleTrees and $\chi$.
+2.3.1. Data Structure With Efficient WM Trade-off
+We start with the STBϵ-tree, observing that its desirable properties derive from the trunk/branch separation, whereas the properties we would like to change derive from their compaction and flush policies.
+We coin the term Bϵ+-tree to describe the class of dictionaries which retain the trunk/branch separation, but which leave open the questions of flush/compaction policy and how (if at all) branches are structured within their referent trunk node(s) (fig.
+2).
+Our first key idea is to instantiate the Bϵ+-tree with a balanced, level-tiered compaction policy which organizes branches into sub-levels within a trunk node, as shown in Figure 5a.
+This data structure we name the TurtleTree.
+2.3.2. System Design With No In-Memory Bottlenecks
+TurtleKV adopts a design similar to LSM-tree-based systems like RocksDB, with a write-ahead log (WAL), in-memory component (MemTable), and log-structured on-disk component (Checkpoint, a TurtleTree).
+The primary bottleneck in this design comes from the $\log(N)$ amplification when merging MemTables into the current Checkpoint.
+We observe that the most CPU-intensive part of this process is the key comparisons for level-to-level merging, and the most memory-bandwidth-intensive part is the copying of key-value data to serialize compacted data at its new location.
+Our second key idea is to use data-parallel algorithms for level merging, and fork/join style multi-threaded leaf page serialization for Checkpoint writes.
+2.3.3. First-Class Write-Memory (WM) Tuning Knob
+Inspired by our RocksDB case study, we identify MemTable size $\overline{M_{w}}$ as our WM tuning knob.
+This defines the distance, in total bytes of updates, between Checkpoints.
+To ensure that increasing checkpoint distance does not degrade in-memory query performance, we build a single large MemTable between checkpoint updates.
+To keep tail latencies under control, and to avoid affecting the on-disk structure of checkpoints, we break finalized MemTables into leaf-page sized batches by doing a key-order scan, applying the batches to the Checkpoint TurtleTree one at a time in-cache until the MemTable is drained.
+This is our third key idea, which we call checkpoint distance tuning with Big MemTables.
+3. TurtleTrees
+In this section we describe the TurtleTree data structure, its core algorithms (batch update, get, and scan), and how it leverages checkpoint distance as a WM trade-off tuning mechanism.
 3.1. TurtleTree Structure
-TurtleTrees are Bϵ+-trees ($\S$2.2) whose per-node update buffer is organized as a fixed-capacity, level-tiered LSM-tree with fanout F=2F=2.333That is, level to level growth factor; not to be confused with the branching factor of the node, or number of pivots, which we denote as $\rho$.
-The structure of TurtleTree node and leaf pages is shown in Figure 4.
+TurtleTrees are Bϵ+-trees ($\S$2.3) whose per-node update buffer is organized into a fixed number of levels as shown in Figure.
+5a.
+The structure of TurtleTree node and leaf pages is shown in Figure.
+5b.
+3.1.1. Leaf Pages
 Leaf pages contain a sorted list of key/value pairs.
-As is typical of B+-trees, leaves are sized several orders of magnitude larger than interior nodes.
-[^4] A typical configuration might be 4KB node pages and 16MB leaf pages.
-To ensure fast in-memory queries (both point and range/scan), TurtleKV leaf pages contain a cache-oblivious, byte-addressed binary trie index.
-This index contains every 16th or 32th key from the leaf, depending on available space.
-The trie is made cache-oblivious by laying out nodes in Van Emde Boas order (Bender et al., 2007) to maximize locality of reference along any given query path.
-Each (interior) node is split into roughly equal halves, with one half devoted to pivots (dark gray) and the other to the update buffer (yellow).
-Within a node page, the update buffer region contains pointers to segments (the blue boxes with rounded sides), which are stored as leaf pages with the addition of tombstone records for deleted keys.
+To achieve high I/O utilization during updates and scans, leaf pages are sized several orders of magnitude larger than interior nodes.
+[^9] The configuration we use is 4KB node pages and 32MB leaf pages.
+In order to guarantee low latencies during point queries, leaf pages are internally structured to allow small slices, or sharded views, to be cached and re-used to find queried keys.
+Leaves also include a cache-oblivious, byte-addressed binary trie index at the front of the page.
+This index contains every $k^{th}$ key from the leaf, $k\in\{1..32\}$.
+The trie is made cache-oblivious by laying out nodes in Van Emde Boas order (benderCacheobliviousStreamingBtrees2007a) to maximize locality of reference along any given query path, minimizing CPU cache misses for in-memory queries.
+3.1.2. Node Pages
+Each (interior) node is split into roughly equal halves, [^10] The intra-page regions in Figure 5b are not shown to scale.
+with one half devoted to pivots (dark gray) and the other to the update buffer (yellow).
+Within a node page, the update buffer region contains pointers to segments, which are stored as leaf pages with the addition of tombstone records for deleted keys.
 The segments in a buffer are stratified into levels of exponentially increasing size.
-Keys are required to be unique and sorted within a level.
+Keys are required to be in-order and unique within a level.
 The first level contains at most a single segment, the second level at most two, the third level four, then eight, etc.
 All levels start out vacant and can be in either a vacant or occupied state.
-For a given number of pivots $\rho$ in a node, the number of levels is set to $\lceil\log_{2}{\rho}\rceil$, and the total number of segments in all levels of the buffer bounded to at most $\rho-1$.
-Figure 4. Details of the internal structure of TurtleKV’s TurtleTree implementation.
+For a given number of pivots $\rho$ in a node, the number of levels is limited to $\lceil\log_{2}{\rho}\rceil$, and the total number of segments in all levels of the buffer limited to at most $\rho-1$.
 Node buffers maintain certain metadata to speed up various operations.
@@
 The default policy is to flush one leaf-sized batch after each batch insert, to the pivot with the highest pendingBytes count, if possible.
-3.2. TurtleTree Batch Updates
+3.2. TurtleTree Operations
+3.2.1. Batch Updates
 TurtleTrees are designed to support efficient batched updates; individual key modifications can be logged/buffered outside the tree and applied as a batch once enough have accumulated.
@@
 The root node is the only node in the tree allowed to have an arbitrarily small number of child pivots, so long as it is at least 2.
-Figure 5. Node-Buffer Insertion of Batch Updates in TurtleTrees.
 To perform a batch update on a node, we first insert the incoming batch into the node’s buffer.
@@
 This cascade of merges produces the final sorted key sequence $\langle 0,1,2,3,4,5,6,7,8,9,10,11\rangle$, resulting in a new level comprised of segments e $\langle 0,1,2\rangle$, f $\langle 3,4,5\rangle$, g $\langle 6,7,8\rangle$, and h $\langle 9,10,11\rangle$.
+Figure 6. Insertion of Batch Updates in TurtleTrees.
 After inserting a new batch into the buffer, if there are at least $L$ (the size of a leaf page) buffered bytes of updates addressed to one of the pivots, then we perform a flush.
@@
 A quotient filter per leaf page is built and flushed when batch updates are finally written to storage.
-3.3. Checkpoint Distance
-Figure 6. The effect of checkpoint distance ($\chi$) on write amplification. (Pages shown are from the example in Figure 5.)
-The $\chi$ parameter controls how frequently newly created pages (nodes, segments, and leaves) are written from memory to storage.
+3.2.2. Queries
+Point and range queries are the same for TurtleTrees as for Bϵ-trees, with the addition of chasing pointers to segment page(s) in the buffer (instead of just reading its contents).
+We use the active pivots bit sets stored for each segment to narrow the search to only those which might contain the key.
+Further pruning of the set of candidate pages to search is done by querying the quotient filter associated with a given segment or leaf.
+After finding a key in a segment, we compare its position in the leaf to the flushed upper bound for its corresponding pivot; if this position is less than the upper bound (i.e., it has been marked as flushed), we act as if the key is not present and continue the search.
+Deletes are performed alongside Put operations as part of a batch update, by setting a key’s value to the special DELETED tombstone value.
+This value is preserved in leaves until the bottom of the tree is reached; at this point, the key is removed from the merged output.
+3.3. TurtleTree Cost Analysis
+3.3.1. Update I/O cost
+Individual key updates travel from root to leaf of the tree as batch updates and flushes take place.
+Since there are $O(\log_{2}{\rho})$ levels within the buffer of each node, and the tree height is $O(\log_{\rho}{(N/L)})$, each key update is written $O(\log_{2}{(N/L)})$ $O(\log_{2}{\rho}\cdot\log_{\rho}{(N/L)})$ times, regardless of page size or branching factor $\rho$.
+[^13] Note that in order to correctly analyze the cost of a key update, we need only consider its initial journey from root-to-leaf, since each time a key’s page is re-written after it reaches the bottom of the tree, we can amortize that cost against the incoming batch, which represents an equal number/size of keys ending their root-to-leaf journey.
+If we define the transfer size $B$ in units equal to a single key update, the I/O cost of a key update is $O(B^{-1}\cdot\log_{2}{(N/L)})$.
+3.3.2. Query cost
+Point queries (Get) differ in their I/O cost depending on whether or not the key is found.
+Here we give a worst-case bound for the general case; in practice this can be dramatically improved through the addition of per-leaf filters, as described in $\S$4.1.2.
+The cost analysis here also depends on the specific I/O model used; if we are counting total page loads, as in the Disk Access Model (DAM), then a single-key (point) query must load all node pages and buffer segments along a single root-to-leaf path: $O(\log_{\rho}{(N/L)}+\frac{L}{B}\cdot\log_{2}{(N/L)})$ or just $O(\frac{L}{B}\cdot\log_{2}{\frac{N}{L}})$.
+Like other B-tree variants, range queries of length $k$ on Turtle Trees cost a point query (to find the starting key) plus $O(k/B)$ IOPs.
+There is also an in-memory, per-key cost of $\log_{2}{(\log_{2}{(N/L)})}$ to merge sorted runs, which is the same as a level-tiered LSM-tree.
+This per-key cost expression is derived via the use of a binary min-heap to select which of the $\log_{2}{(N/L)}$ levels of the tree contains the least-ordered key.
+3.3.3. Write-Memory Trade-off Efficiency
+The WM tuning knob for TurtleTrees is a parameter chi ($\chi$), for checkpoint distance, which determines the number of batches applied to a TurtleTree in-memory before flushing changed pages to storage as a new checkpoint.
 At $\chi=1$, every time we perform a batch update on the tree, all modified pages must be written back out.
@@
 If $\chi$ is doubled again to 4, then all keys in the example appear to skip the first two levels.
-3.4. Cost Analysis & Limitations
-Table 1 compares TurtleTrees to other data structures.
-We see that TurtleTree’s update put (update) operations benefit greatly from checkpoint distance due to the $\chi$ parameter in the denominator.
-Gets under different memory models perform similarly to other data structures, but are greatly increased by read caching (not shown in the table).
-We next describe how the table values are derived.
-Individual key updates travel from root to leaf of the tree as batch updates and flushes take place.
-Since there are $O\left(\log_{2}{\rho}\right)$ levels within the buffer of each node, and the tree height is $O\left(\log_{\rho}{\left(N/L\right)}\right)$, each key update is written $O\left(\log_{2}{\rho}\cdot\log_{\rho}{\left(N/L\right)}\right)=O\left(log_{2}{\left(N/L\right)}\right)$ times, regardless of the node page size or tree branching factor $\rho$.
-Note that in order to correctly analyze the cost of a key update, we need only consider its initial journey from root-to-leaf, since each time a key’s page is re-written after it reaches the bottom of the tree, we can amortize that cost against the incoming batch, which represents an equal number/size of keys ending their root-to-leaf journey.
-If we define the transfer size $B$ in units equal to a single key update, the I/O cost of a key update is
-$O\left(\frac{1}{B}\cdot\log_{2}{\frac{N}{L}}\right)$
-Point queries (Get) differ in their I/O cost depending on whether or not the key is found.
-Here we give a worst-case bound for the general case; in practice this can be dramatically improved through the addition of per-leaf filters.
-The cost analysis here also depends on the specific I/O model used; if we are counting total page loads, as in the Disk Access Model (DAM), then a single-key (point) query must load all node pages and buffer segments along a single root-to-leaf path: $O\left(\log_{\rho}{\frac{N}{L}}+\frac{L}{B}\cdot\log_{2}{\frac{N}{L}}\right)$ or just $O\left(\frac{L}{B}\cdot\log_{2}{\frac{N}{L}}\right)$.
-However, because “latency lags bandwidth” (Patterson, 2004), it is also reasonable to consider I/O models where latency and parallelism are explicitly accounted for (Bender et al., 2021, 2019; Aggarwal and Vitter, 1988), e.g.
-the Parallel Disk Access Model (PDAM).
-Under the PDAM model, Turtle Trees only incur a cost of $O\left(\log_{\rho}{\frac{N}{L}}\right)$ parallel I/Os per point query: one per node on the key’s root-to-leaf path.
-This makes point queries on Turtle Trees on-par with (non-buffered) B-trees.
-Like other B-tree variants, range queries on Turtle Trees cost a point query (to find the starting key) plus $O\left(\frac{k}{B}\right)$ IOPs.
-There is also an in-memory, per-key cost of $\log_{2}{(\log_{2}{(N/L)})}$ to merge sorted runs, which is the same as a level-tiered LSM-tree.
-This per-key cost expression is derived via the use of a binary min-heap to select which of the $\log_{2}{N/L}$ levels of the tree contains the least-ordered key.
-How $\chi$ effects cost analysis. We claim that, in general, the effect of $\chi$ is that keys skip the first $\log_{2}{\chi}$ buffer levels of the tree, which reduces the per-key write cost to:
+Figure 7. The effect of checkpoint distance ($\chi$) on write amplification. (Pages shown are from the example in Figure 6.)
+We claim that, in general, the effect of $\chi$ is that keys skip the first $\log_{2}{\chi}$ buffer levels of the tree, which reduces the per-key write cost to:
 $O\left(\frac{1}{B}\cdot\left(\log_{2}{\frac{N}{L}}-\log_{2}{\chi}\right)\right)=O\left(\frac{1}{B}\cdot\log_{2}{\frac{N}{\chi L}}\right)$
-Limitations.
-The key limitation of TurtleTrees and $\chi$ is memory usage to support $\chi$ instead of read caching.
-But, $\chi$ can be tuned dynamically during runtime so $\chi$ need only be increased when needed.
-In return, TurtleTrees and $\chi$ pay less read-filter cost for the lifetime of data.
-4. From TurtleTrees to TurtleKV
-Figure 7 shows TurtleKVs high-level architecture.
-It is similar to systems based on log-structured B-trees and on LSM-trees (Figure 7).
-Per-key updates are inserted into an in-memory index, or MemTable, and appended to a write-ahead log (WAL) for durability.
-The purpose of the MemTable, as in LSM-tree designs, is to buffer updates until we have reached a minimum batch size.
-Once this happens, the checkpoint pipeline sorts and de-duplicate key/value pairs in the MemTable, producing a batch which is applied to the most recent checkpoint (which is a TurtleTree).
-This process is repeated until $\chi$ batches (MemTables) have been applied, at which point the modified portions of the TurtleTreeare serialized and written to storage.
-Afterward, the WAL is trimmed.
-Queries are processed by first consulting the MemTable, which represents recent updates (up to some size threshold).
-If the key is found here, the query finishes.
-Otherwise, the query proceeds to check recent MemTables in most-recent order, one per batch applied to the in-memory pending checkpoint.
-If the key is not found in any of these, the most recent durable checkpoint is searched.
-Implementation & parameters We implemented TurtleKV in approximately 30k lines of C++ code.
-TurtleKV’s static parameters include $B$ (the block size), $M$ the total memory budget, the filter bit rate ($FBR$), and the number of threads involved in various stages of updating Memtables and the checkpoint pipeline ($T$).
-In this work, we investigate the value of dynamically changing checkpoint distance ($\chi$) on demand, This parameter varies the delay between creating pages or batches and writing them to disk.
-TurtleKV needs to address two high-level challenges.
-First, it must guarantee high throughput and low latency.
-Second, for $\chi$ to have value, it must guarantee that I/O, not memory or CPU is the bottleneck.
-We next describe TurtleKV’s key components in more detail and how we address these challenges for them.
-Figure 7. High-level Architecture of TurtleKV
-4.1. TurtleKV Component Details
-Write-Ahead Log.
-TurtleKV writes all key updates to a shared change log file initially, to provide data durability in the event of program termination before the next checkpoint update.
-This file is divided into fixed-sized blocks, which internally use a slotted page format to store variable-sized per-key edits.
-To avoid low performance due to contention when logging updates, we implement a separation-of-concerns strategy where each user thread writes key updates to a thread-local block buffer to minimize contention.
-A single background thread is responsible for flushing data to storage.
-The background flusher thread uses a lock-free, periodic polling scheme to steal ownership of buffers from foreground threads.
-The polling period can be adjusted to trade off between minimizing the latency for updates to become durable, and the throughput achieved by update batching.
-Change log blocks may be configured large enough (e.g., ¿ 4kib) so that writing to storage is not atomic.
-Therefore the block header includes a hash of its contents.
-At recovery time, any block whose hash does not match is assumed to be partially written, and is ignored.
-Each update to the change log is assigned a unique, monotonic sequence number using an atomic counter.
-This can be used during recovery to maintain a consistent order of events across restarts.
-MemTables.
-We use Adaptive Radix Trees (ARTs) as the core data structure for indexing buffered updates (Leis et al., 2013).
-This allows TurtleKV to maintain an ordered index for range queries, providing low latency.
-A concurrent hash table index can also be enabled at runtime to further speed point queries, at the cost of slightly higher insertions since the ART is still used.
-To mitigate contention between multiple writers, TurtleKV uses optimistic lock coupling (OLC) (Leis et al., 2016).
-We also exploit the fact that MemTables do not need to support true key removals, since deletions must be merged into the checkpoint tree.
-This allows us to use a simple arena-based memory allocation scheme, with each thread allocating memory from the system in 64kib chunks which are then divided as needed into ART nodes.
-TurtleKV MemTables rely on WAL block buffers to store actual key/value data, further simplifying the job of the ART index, since it does not need to copy update data.
-Checkpoint Pipeline.
-We use the parallelism of the underlying hardware to perform the needed work.
-A foreground thread first places finalized Memtables in a queue, which is consumed by a background thread.
-This thread sorts and compacts key/value pairs to produce new batches of updates.
-Each batch is placed in another queue which is consumed by the thread responsible for updating the checkpoint tree.
-The checkpoint pipeline must also manage copying of data as sorted key/value pairs are merged/compacted within buffer levels and flushed down the tree.
-By increasing $\chi$, we expect incoming key updates to be pushed further down the tree before they are written to storage.
-If we naïvely copy key/value data into page buffers during this process at every step, most of this work will be wasted in building pages which are never written.
-To address this challenge, Checkpoint tree updates defer the copying of key/value data until the last batch has been applied (according to the $\chi$ parameter).
-Merge/compactions on in-memory checkpoint tree update buffer levels and leaves therefore operate on pointers, to avoid data copies.
-The individual merge/compact operations are parallelized using a background thread pool.
-This allows write-heavy workloads that execute on only a few threads to take advantage of available CPU cores when possible.
-When the final batch for a checkpoint has been applied, all new pages (leaf, node, and filter) are built concurrently using the same thread pool used for merging.
-Once all pages have been built, the checkpoint is handed via a third queue to the final thread, which writes the page data in parallel to storage.
-Page lifetime is controlled via reference counting.
-The reference counts, as well as root references to subsequent checkpoint trees, are appended to a write-ahead log using a two-phase commit protocol to ensure crash consistency.
-Once a new checkpoint has been made durable, the checkpoint field of the KVStore is updated and the finalized MemTables representing the batches used to create the new checkpoint are pruned from the Deltas stack.
-Pruning the MemTables from Deltas allows the change log to be trimmed, releasing storage resources and ensuring that recovery times do not grow unbounded.
-Page Cache.
-TurtleKV uses LLFS (Astolfi et al., 2024), MathWorks’ Low-Level File System library.
-This means that storage resources are organized into different pools (extents) of storage, one per page size.
-[^7] Typically one for nodes (4kb), one for leaves/segments (2mb-16mb), and one for filters (16kb-256kb).
-Each of these pools is represented in LLFS as a PageDevice.
-While it is advantageous to use large (>2>2mb) page sizes for leaves when writing new checkpoint pages, reading entire leaf pages while processing a point query can introduce a severe latency penalty.
-To address this issue, we add a kind of virtual PageDevice to LLFS, called a sharded view.
-Sharded views assign page identifiers to slices of larger pages; e.g., a 4kb sharded view of a 16mb page device allows the loading/caching of 4kb aligned slices.
-This feature allows the point query code path to load and cache only the parts of a leaf page that are needed to answer a query.
-Using sharded views allows the TurtleKV’s large leaf page design to perform competitively in low cache configurations when compared to the alternate Bϵ+-tree design in SplinterDB, which uses 24mb B+-trees composed of 4kb pages instead of a single large page.
-TurtleKV’s sharded page view design has the additional advantage of being able to answer key presence queries (return true or false depending on whether the key is present, instead of returning its value) more efficiently, since key and value data are stored in separate parts of a leaf.
-Caching policy.
-Even if we optimize all the other steps so that write I/O is the bottleneck, we still need to achieve a high cache hit-rate when updating the TurtleTree checkpoint.
-This is because the I/O improvements we obtain from increasing $\chi$ may be offset by increased I/O from reading pages containing data to merge, compact, or flush.
-To address this issue, we use use a modified CLOCK (Jiang et al., 2005) cache replacement policy where each page has a priority value (instead of a single bit) which is decremented when the clock hand passes it.
-This allows TurtleKV to provide information to the lower layer about which pages to make an extra effort to keep in memory.
+The key limitation of TurtleTree’s $\chi$ tuning knob, and of RM+WM tuning schemes in general, is that unlike I/O operations, memory is generally a bounded quantity.
+This means higher $\chi$ value will spend memory that could have been used for page and filter caching.
+Thus the zero-sum of fixed-size memory resources is effectively turned into a read-vs-write (RW) trade-off.
+However, unlike RW trade-offs that affect the structure of stored data, our scheme is flexible and forgiving, since it can be changed at runtime.
+4. TurtleKV System Design
+In this section we first give an overview of the system architecture and update/query paths in TurtleKV, then dive deeper into the design of two of the key ideas it implements: checkpoint distance tuning with Big MemTables, and parallel (multi-core) acceleration for TurtleTree (checkpoint) updates.
+4.1. Basic Operation
+TurtleKV is based on a conventional LSM-tree system architecture employing three high-level components: a write-ahead log (WAL) to ensure durability of incoming updates, an in-memory index and buffer of recent data (MemTable), and an on-disk checkpoint for long-term storage (TurtleTree).
+It is implemented in  32k lines of C++ code, and is available as open-sourced software (HttpsGithubcomMathworks2025).
+4.1.1. Update Path
+Incoming key/value data are copied to an internal memory buffer which is continuously flushed in the background to the WAL.
+Each update is assigned a unique sequence number so that the database state can be consistently reconstructed during recovery.
+After the key and value bytes are copied, they are inserted into an in-memory index (the active MemTable) on the user thread performing the update; as soon as this completes, the update is visible to all threads in the process.
+Each MemTable is limited to a maximum size $\overline{M_{w}}$.
+When a thread attempts to perform a MemTable insertion, but fails because this limit would be exceeded, the active MemTable is finalized, or made read-only, and a new active MemTable is created to accept future updates.
+The finalized MemTable is handed off to a background thread, which splits it into 1 or more leaf-page-sized batches, each of which is applied to the current checkpoint TurtleTree.
+The user thread must wait if the background thread is already busy processing a previous finalized MemTable.
+After the hand-off succeeds, it will also wait until the total number of finalized MemTables in the system is at most 2.
+When the background thread finishes applying all batches for a MemTable to the page cache, it hands the list of new pages to a second background thread (the writer thread), which writes the new checkpoint data to disk.
+Once the new TurtleTree pages are durable, the writer thread swaps the finalized MemTable with a reference to the new checkpoint which subsumes it, possibly unblocking any application threads that were waiting for the finalized MemTable count to decrease.
+This multi-stage update pipeline allows high system resource utilization by allowing the different stages (MemTable insertion, TurtleTree updates, and checkpoint page flushing) to be performed at the same time.
+By keeping the amount of queued data in the pipeline to a minimum, it also provides back-pressure so that no part of the system outpaces the others, in order to avoid unbounded tail latencies.
+4.1.2. Query Path
+Point queries are answered by first consulting the active MemTable.
+If the target key is found there, the query terminates and the answer is returned to the caller.
+Otherwise, the $\leq$2 finalized MemTables are consulted in newest-to-oldest order.
+If the key is not found there either, then the most recent checkpoint is queried as described in $\S$3.2.2.
+TurtleKV supports either Bloom or Quotient filters (selected at compile-time) to optimize point queries.
+Each leaf page, whether at the bottom of the TurtleTree or referenced from a node buffer, has an associated filter page.
+Whenever the query algorithm demands that a leaf be searched for a given key, the associated filter is first queried, and only if it gives a positive response is the leaf loaded/checked.
+When a leaf is queried, we first attempt to pin the whole leaf page in the cache; this will fail if the page isn’t in memory, without performing any I/O.
+If the whole leaf is not cached, then we access the first  64kb-sized slice of the page, which contains the header and trie index.
+From this information, we are able to determine which 4kb slice will contain the target key, if present.
+4.2. Parallelization of TurtleTree Updates
+4.2.1. The necessity of parallelized updates
+TurtleTree checkpoint update is the most expensive operation in the TurtleKV update path by a substantial margin.
+Therefore it is the primary target of our optimization efforts to avoid in-memory bottlenecks, which is one of the 3 high-level requirements for TurtleKV ($\S$2, Table 1).
+[…]
+The challenge of TurtleTree updates is that they represent a potentially logarithmic amplification of work as compared to earlier pipeline stages, due to the $\log(N/L)$ levels through which update records must be merged/compacted to reach the leaf level.
+There is simply no way to keep up with load generated by earlier stages of the update path if we do not employ some kind of in-memory parallelization strategy.
+4.2.2. Task-parallel vs Data-parallel
+There are two basic approaches to this: allow concurrent updates to take place in different locations within the same tree, or somehow parallelize the work of sequentially applied batch updates.
+The former approach (task-parallel) is used by SplinterDB; TurtleKV takes the latter approach (data-parallel).
+The reason is twofold: first, there are many complications involved in implementing concurrent updates to Bϵ+-trees, and these are a superset of the complexities one faces when implementing concurrent B+-trees, a non-trivial task because of the competing concerns of root-to-leaf traversal for insertion, and leaf-to-root split/merge for re-balancing.
+The second reason is, as we explain below, that adopting a data-parallel approach allows us to better control load balancing across processors, leading to more consistent update performance with less variability.
+4.2.3. Details of parallel updates
+We observe that the most CPU-intensive operations in TurtleTree batch update are the key comparisons required to merge/compact level segments when a node receives an incoming batch from its parent and flushes an outgoing batch to a subtree.
+Merging multiple ordered sequences is easily done in parallel using multi-selection (deoOptimalParallelAlgorithm1994) on each sequence to determine the input and output ranges for different processors.
+By implementing this algorithm, we are able to scale the CPU work to multiple processors while ensuring the correctness of our implementation through simplicity of the code.
+We also observe that the most memory-bandwidth expensive operation during checkpoint update is copying key-value data while serializing new compacted pages prior to writing them to storage.
+Our approach to optimizing this is reduce the total amount of data that must be copied by deferring all serialization until right before the checkpoint is finalized, and to serialize all leaf pages (and their corresponding filters) in parallel on different threads.
+4.3. Checkpoint Distance & Big MemTables
+4.3.1. Preliminaries: MemTable Implementation
+The requirements for the MemTable data structure are scalability to high numbers of concurrent threads, low space overhead, ability to perform key-order scans, and CPU cache friendliness.
+There are several possible data structures which meet these criteria; for example, Wang et al.
+(wangBuildingBwTreeTakes2018) compare Bw-Tree, OpenBW-Tree, SkipList, Masstree, B+-tree, and Adaptive Radix Tree (ART) implementations for in-memory, multi-threaded performance and efficiency.
+TurtleKV implements ARTs for this purpose, because they outperform all other data structures in this comparison, and offer low memory usage and high CPU-cache hit rate across a variety of workloads.
+However, TurtleKV is careful to abstract this implementation choice so that other data structures can be swapped in if desired.
+For comparison, RocksDB uses concurrent SkipLists as the default MemTable, WiredTiger uses a combination of B+-tree and SkipLists for in-memory indexing, and SplinterDB uses concurrent B+-trees as its MemTable and branches.
+The TurtleKV ART implementation is roughly the same size (lines of code) as RocksDB’s SkipList.
+(a) KOps/Second, 16gb cache
+(b) KOps/Second, 16gb cache
+(c) KOps/Second, 48gb cache
+(d) KOps/Second, 48gb cache
+[…]
+4.3.2. Checkpoint Distance: A Design Dilemma
+TurtleKV’s write-memory (WM) tuning knob is checkpoint distance; this is the total size of the updates between two checkpoints.
+This can be varied at runtime to target different levels of write-optimization at the expense of higher memory footprint devoted to updates.
+Recall from our case study ($\S$2.2.2), that we find there are two knobs offered by RocksDB to trade write amplification directly against memory: MemTable size $\overline{M_{w}}$ and MemTable count $|M_{w}|$.
+The theoretical reason behind their effectiveness is that they force the first $O(\log_{F}(M))$ compactions for a given key $x$ to happen in memory, before that key is ever written to an SSTable.
+Varying either parameter while holding the other constant is equally effective, since $M=\overline{M_{w}}\cdot|M_{w}|$.
+The reason to prefer increasing $\overline{M_{w}}$ to $|M_{w}|$ as a tuning knob is that MemTables’ contribution to query cost in a RocksDB-like design is $O(|M_{w}|\cdot\log(\overline{M_{w}}))$, assuming a MemTable data structure that supports optimal logarithmic query cost.
+If we apply this same logic to TurtleKV’s design, however, we encounter a dilemma.
+We want the on-disk structure of checkpoint TurtleTrees to be independent of checkpoint distance, so that we get a true WM tuning knob; this is possible only if we keep $\overline{M_{w}}$ constant and vary the number of batches $|M_{w}|$ in between checkpoints.
+Thus we have a tug-of-war between the concerns of in-memory (MemTable) queries on newer data, and queries against checkpoints (TurtleTrees) in the face of varied past tuning point selection.
+4.3.3. The Solution: Big MemTables
+We resolve this tug-of-war by setting the size limit of a MemTable when it is created to the current checkpoint distance, and breaking up finalized MemTables into many leaf-page sized batches in order to update the TurtleTree.
+We do this via a least-to-greatest key-ordered scan of the MemTable.
+This design is called Big MemTables in TurtleKV.
+Big MemTables help to optimize TurtleKV’s performance in multiple ways.
+First, by scaling up the size of TurtleKV’s in-memory index on recent updates, we speed up queries on ”hot keys.” Second, by allowing a larger set of updates to accumulate in memory before doing a scan to break the MemTable into batches, we increase the skew in key distribution when applying batches to the checkpoint.
+TurtleTree batch update ($\S$3.2.1), which flushes batches down the tree as soon as the minimum size threshold for a subtree is reached (not when the buffer fills up), transparently leverages this key skew to move data down the tree faster, so long as doing so will not hurt efficiency (i.e., so long as we preserve the property that merges are always performed on equal-sized inputs).
+We note that SplinterDB’s flush-then-compact policy achieves a similar effect, though it has the downside of increased space amplification due to the fact that a flushed branch may end up referenced in many different subtrees despite only containing a small amount of live data for each reference.
+Finally, Big MemTables further motivate our selection of ARTs as MemTable data structure.
+Because the structure of radix trees allow for segments of indexed keys to be shared, it is possible for the per-key overhead of ARTs to fall below the average key length (leisAdaptiveRadixTree2013); this efficiency is often increased as the total set of keys in an ART scales.
+Thus by increasing $\overline{M_{w}}$ and then splitting buffered updates into key-skewed batches, we not only play to the strengths of TurtleTrees, but to those of ARTs as well.
 5. Evaluation
-We claim that TurtleKV demonstrates how a key-value storage engine can be built to support low cost re-tuning at runtime, with a dynamic trade-off range that provides excellent resource utilization and competitive performance for write-heavy, read-heavy and mixed workloads.
-The goal of the experiments in this section is to support this claim.
-We aim to answer the following questions:
-  • Question 1: Does TurtleKV offer competitive throughput, write amplification, and space amplification for write-heavy workloads?
-  • Question 2: Does TurtleKV offer superior latency and throughput in read-heavy workloads, dominated by point and range queries?
-  • Question 3: What are the properties, both in cost and benefit, of runtime adjustments to checkpoint distance to select an optimal tuning point in response to workload?
-Methodology.
-To answer these questions, we use the Yahoo Cloud Serving Benchmark (YCSB), comparing TurtleKV to other state-of-the-art key-value engines, each representing a different optimization target: RocksDB, a write-optimized and widely used LSM-tree-based system; and SplinterDB, a state-of-the-art hybrid-optimized key-value store.
-Splinter and Rocks are run mostly in their default configurations; direct I/O was enabled in RockDB in order to control the size of the block cache (this resulting in a slight improvement in overall throughput).
-In our benchmarking, we use two record sizes, 128 bytes and 1024 bytes, 8 bytes of which is the key.
-Each workload is preceded by an initial load phase, where we populate the database with 48GiB of data, 400M 128-byte records or 50M 1024-byte records.
-We test both out-of-cache and in-cache workloads; for out-of-cache, we limit cache size to 16GiB, $1/3$ of the pre-loaded data size, and for in-cache we allow up to 48GiB of cached pages.
-In our parameter selection process, we found that for out-of-cache workloads, a relatively small number of threads is needed to reach the point where I/O becomes the bottleneck.
-Correspondingly, for in-cache, the bottleneck shifts to CPU and memory bandwidth.
-Consequently, our out-of-cache tests use 4 client workload threads, and in-cache tests use 16 threads.
-We also include results for single-threaded tests to show latency of sequentially issued requests.
-Hardware setup.
+In this section, we explore the end-to-end performance of TurtleKV, comparing it to WiredTiger, RocksDB, and SplinterDB.
+We provide data to support the claims we make about TurtleKV and its novel data structure TurtleTrees, and our claims about the necessity of write-memory (WM) trade-off tuning more generally.
+(a) KOps/Second, 16gb cache
+(b) KOps/Second, 16gb cache
+(c) KOps/Second, 48gb cache
+(d) KOps/Second, 48gb cache
+(e) WAF vs Checkpoint Dist. (MB)
+Figure 9. TurtleKV WM Tuning Sensitivity
+5.1. Experimental Setup
+5.1.1. Hardware setup
 Experiments were run on an AMD Ryzen Threadripper 7970x workstation (32 cores/64 threads), with 128 GB of DDR5-6000 registered ECC memory.
@@
 The CPU, while physically packaged as a single-socket part, was configured in the firmware (ASRock TRX50WS) to report its topology as a 4-node NUMA processor, one node per chiplet, each with 8 cores/16 threads.
-(a) 400M $\times$ 128B, 1 thread
-(b) 400M $\times$ 128B, 4 threads
-(c) 400M $\times$ 128B, 16 threads
-(d) 50M $\times$ 1024B, 1 thread
-(e) 50M $\times$ 1024B, 4 threads
-(f) 50M $\times$ 1024B, 16 threads
-Figure 8.
-YCSB Results (higher is better): TurtleKV matches or exceeds performance of other engines on small records, and is competitive on large records.
-For 1024B records, TurtleKV performs better in higher cache configs due to separation of key and value data in large (16mib) leaf pages degrading query performance under cache pressure.
-(a) 400M $\times$ 128B, Write Amp
-(b) 50M $\times$ 1024B, Write Amp
-(c) 400M $\times$ 128B, Space Amp
-(d) 50M $\times$ 1024B, Space Amp
-Figure 9. Write & Space Amplification (lower is better): TurtleKV has slightly higher WAF than SplinterDB, but uses less than half the space on disk.
-(a) 400M $\times$ 128B, Workload E (scans)
-(b) 50M $\times$ 1024B, Workload E (scans)
-Figure 10. TurtleKV’s balanced compaction policy allows it to achieve high throughput when performing short range queries (<<100 records).
+5.1.2. Software/Benchmark Configuration
+We use the Yahoo Cloud Serving Benchmark (YCSB)(cooperBenchmarkingCloudServing2010), comparing TurtleKV to other state-of-the-art key-value engines, each representing a different optimization target: WiredTiger, a read-optimized B+-tree-based engine; RocksDB, a write-optimized and widely used LSM-tree-based system; and SplinterDB, a state-of-the-art hybrid-optimized key-value store.
+Each YCSB run inserts 400M records comprised of 8 byte keys and 120 byte values (Load), then performs 150M operations comprised of 50% updates/50% gets (workload A), another 150M operations comprised of 5% updates/95% gets (workload B), 150M gets (workload C), 15M operations comprised of 95% scans (len=1..100)/5% updates (workload E), and then a final 150M operations comprised of 50% read-modify-writes/50% gets (workload F).
+The total size of the initial data set is 48gb; to evaluate both out-of-cache and in-cache performance, we run the benchmark twice for each database, once with cache size 16gb ($\frac{1}{3}$ of data size) and again with 48gb cache (large enough to fit the raw data).
+5.1.3. Database Configurations
+The configuration used for each database is informed by the case studies described in $\S$2.2.
+For WiredTiger, we give a generous allocation to write buffering in cache, setting eviction dirty target to half the cache size and eviction dirty trigger to 95% of the cache.
+We turn on transaction-level logging for durability, and disable automatic checkpointing for our tests.
+For RocksDB, we set parallelism equal to the number of cores on the test machine (64); we also set the maximum number of parallel sub-compactions to this number.
+We set the write buffer size equal to 64mb, since in our testing we did not observe an end-to-end advantage to larger values (despite lower write amplification).
+We enable the block cache explicitly in order to accurately track the memory RocksDB uses for page caching (otherwise it uses the kernel file cache), and also enable direct I/O, which we observe to be slightly faster.
+SplinterDB is configured with mostly default values.
+We set the maximum key size to 8 bytes to match our test workload, and configure the number of background threads to 2 for memtable tasks, 24 for normal tasks, following the guidance of the SplinterDB tuning guide.
+(a) Insert, (16gb)
+(b) Update (16gb)
+(c) Get (16gb)
+(d) Scan (16gb)
+(e) Insert (48gb)
+(f) Update (48gb)
+(g) Get (48gb)
+(h) Scan (48gb)
+Figure 10. YCSB, Per-Operation Latency Distributions for 16gb and 48gb Cache Sizes
 […]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Figure 11.
-Checkpoint distance ($\chi$) tuning reduces writes and put latencies (a, e), increases upsert throughput, and increases memory usage (b, f).
-Adjusting $\chi$ at runtime is transparent (c, g) and cheap, costing only a few seconds to flush the current checkpoint (d, h).
-5.1. Experiments
-To evaluate write performance, we run YCSB Load (insertion of unsorted records into an empty database), Workload A (50%/50% mixture of reads and writes), and Workload F (50% reads and 50% read/modify/write).
-We measure and compare throughput (operations/sec), write amplification, space amplification, and memory usage.
-Generally, we are looking for high insert/update throughput (with correspondingly low average latency), low write and space amplification, and peak memory usage generally in line with the configured cache size settings.
-To evaluate read performance, we run YCSB Workloads B (95% reads, 5% updates), C (100% point queries), and E (95% scans from 1 to 100 keys in length).
-We expect that the level-tiered structure of TurtleTrees will provide a performance boost in read-heavy workloads.
-We explore the dynamic tuning capabilities of TurtleKV via a modification to the YCSB-C benchmark implementation.
-Before and after each workload, database drivers are given access to the workload properties and allowed to modify their dynamic configuration in response.
-This mechanism, while not realistic per se, allows an investigation into the possibilities offered by on-demand cache size and $\chi$ adjustment.
-We leave the problem of automatically detecting and adapting to workload characteristics using $\chi$ tuning for future research, for the moment focusing on the properties of the tuning knob itself.
-To evaluate the effect of $\chi$-tuning, we run the YCSB Load with various values of $\chi$, from 1 to the memory limit of the test machine, followed by workloads C (100% gets) and E (95%/5% short scans/puts).
-The $\chi$ parameter is reset to 1 after the initial load and before the read workloads.
-This causes the checkpoint update pipeline ($\S$4.1) to be flushed, writing a new checkpoint to disk, trimming the WAL, and dropping old MemTables.
-We look to demonstrate that write amplification scales proportionally to put latency through the entire dynamic range, for large and very small (64 byte) records, and that the theoretically derived cost reduction of $O(-\log{\chi})$ is observed in practice.
-We also aim to show that the latency of re-tuning, i.e.
-of flushing the last checkpoint, is relatively minor and scales with $\chi$.
-5.2. Results
-Overall we found our experimental data supports the claim that TurtleKV offers performance matching or exceeding state-of-the-art key-value engines, with the flexibility of dynamic tunability.
-Below, we present the key results of our evaluation.
-[…]
-TurtleKV uses less than half of the on-disk space that SplinterDB does for the same data.
-Turtle’s write throughput trails Splinter’s in some cases because it has higher write amplification.
-How much checkpoint distance could reduce this was limited by the available memory.
-If the overhead of maintaining in-memory, unflushed checkpoints can be reduced, our approach might be able to surpass the highly write-optimized, lazily compacted design of Splinter.
-This potential warrants further research.
-[…]
-Result 3: Dynamic checkpoint distance based tuning offers dramatic improvements in write performance, but its memory cost yields diminishing returns in the limit.
-Our data supports the theoretical analysis of the benefits of $\chi$-based tuning.
-Tuning $\chi$ to the limit of available memory allows write amplification to be reduced from 12-14 down to 2-4, a dramatic improvement (Figure 11a, e).
-As claimed, $\chi$-based write optimization carries no residual penalty (Figure 11c, g), as shown by the stable query throughput curves overlaid by increasing insert throughput.
-The re-tuning cost (latency) of $\chi$, which is dominated by the flushing of a partial checkpoint, is under 10 seconds for most values, though it does increase with checkpoint distance (Figure 11d, h).
-The limitation of $\chi$ is that it does offer diminishing returns at higher memory usage levels (Figure 11b, f).
-This presents an opportunity for future research.
+5.2. Results & Discussion
+5.2.1. YCSB Throughput: TurtleKV vs Others
+Figure 8 shows the end-to-end throughput of each YCSB workload, for dynamically-tuned TurtleKV and all other databases.
+For cache-constrained write-heavy workloads (8a), SplinterDB was the highest performer, which agrees with our case studies on WM tuning (SS2.2).
+In the cache-constrained read-heavy workloads (8b), TurtleKV was able to match or exceed the performance of the other write-optimized systems (RocksDB and SplinterDB) in point queries, which is expected because TurtleTree’s level-tiered compaction policy is more read-optimized than the lazy compaction of SplinterDB.
+TurtleKV’s performance in workload E (scans) shows room for improvement; we believe the low scan rate is due to cache pressure from the active (unflushed/uncompacted) MemTable left over from previous workloads.
+This data supports the idea that TurtleKV’s scan performance in cache-limited configurations could be improved by implementing a mechanism to detect query-driven cache pressure and respond by flushing the current MemTable to the checkpoint TurtleTree to free up memory.
+WiredTiger’s B+-tree is the clear winner for workload E in both cache configurations, because it doesn’t have any in-memory merging to do, unlike the other systems.
+TurtleKV performed excellently in the high-memory configuration across all workloads (Figures 8c, 8d), only being bested in scan-dominated workload E again by WiredTiger’s optimal B+-tree.
+5.2.2. Static vs Dynamic Tuning
+Figure 9 (a-d) shows TurtleKV’s sensitivity to checkpoint distance tuning parameter selection across YCSB workloads.
+Particularly noteworthy is the benefit of setting a lower checkpoint distance for point-query dominated workloads B and C (9b, 9d).
+The ability to give explicit direction to the system as to how it should use memory helps TurtleKV to achieve much higher performance in these workloads than the other systems (8b, 8d), including TurtleKV itself when tuned statically.
+Figure 9e demonstrates our claim that the write-optimization benefits of WM tuning in TurtleTrees scales independently of total dictionary size $N$.
+The higher variation in the smallest dataset (N=31M) is because this scenario is more sensitive to small differences in the number of batches flushed down the tree, since the leaf page size is fixed at 32mb, which is a larger percentage of the overall data size at smaller scale.
+5.2.3. YCSB Latency Percentiles
+Figure 10 shows the per-operation latency distributions for the YCSB tests.
+A key result in this data is that TurtleKV demonstrates superior tail latencies (P99.9) for both inserts and updates, which is evidence supporting the effectiveness of TurtleKV’s update pipeline and its multi-processor approach to checkpoint updates.
+The only other system which performed consistently better than TurtleKV in terms of tail latencies is SplinterDB’s point queries.
+These are helped by its use of Quotient Maplets, a state-of-the-art mechanism which could in principle be adapted for use in TurtleTrees.
+(a) 16gb Cache
+(b) 48gb Cache
+Figure 11. Memory Usage over Time During YCSB Benchmarks
+5.2.4. Memory Usage over Time
+Figure 11 shows memory usage in the benchmark process for each database over the course of the YCSB workloads.
+TurtleKV and SplinterDB both quickly make use of available memory during the Load; afterwards, both generally stayed below their allocated limit.
+Note, because the load generator is in the same process as the key-value stores being tested, and because we used pre-recorded workload trace files that are buffered in memory, the overall memory footprint is somewhat higher than the configured cache limit.
+Both RocksDB and WiredTiger did not make use of all available headroom during write-heavy workloads.
+RocksDB did slightly better than TurtleKV and SplinterDB at staying within the configured limit; WiredTiger’s memory overhead caused it to use the most memory over the course of the benchmark.
+Figure 12. CPU Cache Miss Rate, 48gb cache
+5.2.5. CPU Efficiency for In-Cache
+Figure 12 shows the CPU cache miss rate for each workload of YCSB.
+TurtleKV is particularly cache-friendly in query-heavy workloads C and E, due to its careful attention to page layout (cache-oblivious trie indexes at the head of each leaf page) and use of ARTs.
+In mixed workloads A and F, TurtleKV performs similarly in this metric to leaders RocksDB and WiredTiger.
+For the Load workload, CPU cache miss rate is elevated due to the shuffling effect of data-parallel batch compactions while updating the TurtleTree checkpoint; this is on-par with SplinterDB’s cache miss rate for the Load phase.
 6. Related Work
-The theoretical trade-off space constraining external memory indexing is characterized by Brodal and Fagerberg (Brodal and Fagerberg, 2003) for comparison-based dictionaries, and by Conway et al.
-(Conway et al., 2018) for hash-based dictionaries.
-The key insights of these works are echoed and elaborated in practical terms by the RUM Conjecture (Athanassoulis et al., 2016).
-Recent efforts to explore trade-offs in LSM-tree based systems include Monkey (Dayan et al., 2017), a key-value store which optimizes queries through an analytically optimal allocation of memory to Bloom filters at various levels of the tree; Dostoevsky (Dayan and Idreos, 2018), which further improves space-time trade-offs for LSM-based systems by varying compaction policy from lazy at lower levels to greedy at higher ones; Wacky (Dayan and Idreos, 2019), which introduces adaptive level sizing and finer-grained control over compaction policy to even better optimize the trade-off curve; and Cosine (Chatterjee et al., 2021), which builds on the previous three and uses machine learning to optimize data structure design for specific cloud-based key/value storage workloads.
+The theoretical trade-off space constraining external memory indexing is characterized by Brodal and Fagerberg (brodalLowerBoundsExternal2003) for comparison-based dictionaries, and by Conway et al.
+(conwayOptimalHashingExternal2018) for hash-based dictionaries.
+The key insights of these works are echoed and elaborated in practical terms by the RUM Conjecture (athanassoulisDesigningAccessMethods2016).
+Recent efforts to explore trade-offs in LSM-tree based systems include Monkey (dayanMonkeyOptimalNavigable2017), a key-value store which optimizes queries through an analytically optimal allocation of memory to Bloom filters at various levels of the tree; Dostoevsky (dayanDostoevskyBetterSpaceTime2018), which further improves space-time trade-offs for LSM-based systems by varying compaction policy from lazy at lower levels to greedy at higher ones; Wacky (dayanLogStructuredMergeBushWacky2019), which introduces adaptive level sizing and finer-grained control over compaction policy to even better optimize the trade-off curve; and Cosine (chatterjeeCosineCloudcostOptimized2021), which builds on the previous three and uses machine learning to optimize data structure design for specific cloud-based key/value storage workloads.
 We note that Dostoevsky and Wacky in particular share some thematic similarity to TurtleKV’s scheme of optimizing writes without harming future query performance by deferring on-disk compaction of recent updates and compacting older data more aggressively.
-Endure (Huynh et al., 2022) tunes size ratio, memory allocations, and compaction policy to maximize throughput under workload uncertainty.
-Notable work in extending the trade-off tuning potential of B-tree based stores includes the Bw-Tree (Levandoski et al., 2013b; Levandoski and Sengupta, [n.d.]; Wang et al., 2018), a log-structured B-Tree which uses lock-free techniques to achieve low overheads for in-memory operations in high concurrency settings; Bf-Trees, a similar approach which introduces the innovation of caching mini-pages using a circular buffer to reduce I/O amplification while minimizing the overhead of cache maintenance; and the Bwe-Tree (Wang et al., 2024), which describes many practical improvements to the Bw-Tree design to optimize for fast SSDs.
-LSB-Trees (Kim and Lee, 2015) propose an alternate mechanism for storing B-Tree deltas using a combination of physical and logical logging.
+Endure (huynhvldb2022) tunes size ratio, memory allocations, and compaction policy to maximize throughput under workload uncertainty.
+Notable work in extending the trade-off tuning potential of B-tree based stores includes the Bw-Tree (levandoskiBwTreeBtreeNew2013; levandoskiBWTreeLatchFreeBTree; wangBuildingBwTreeTakes2018), a log-structured B-Tree which uses lock-free techniques to achieve low overheads for in-memory operations in high concurrency settings; Bf-Trees, a similar approach which introduces the innovation of caching mini-pages using a circular buffer to reduce I/O amplification while minimizing the overhead of cache maintenance; and the Bwe-Tree (wangBweTreeEvolution2024), which describes many practical improvements to the Bw-Tree design to optimize for fast SSDs.
+LSB-Trees (kimLSBTreeLogstructuredBTree2015) propose an alternate mechanism for storing B-Tree deltas using a combination of physical and logical logging.
 The relationship of Log-Structured B-Trees to the overall continuum of trade-offs in the external memory dictionary design space is explored by Idreos et al.
-(Idreos et al., 2019).
+(idreosDesignContinuumsPath2019).
 Several recent works address key/value engine design in the face of very fast modern storage devices.
-SplinterDB (Conway et al., 2020) and Mapped SplinterDB (Conway et al., 2023) introduce innovations in the design of external memory dictionaries, extending Bϵ-Trees to STBϵ-Trees by applying ideas from size-tiered LSM-trees to achieve high update efficiency without excessively penalizing point queries.
-FASTER (Chandramouli et al., 2018a, b) combines log-structured storage with in-memory concurrent hash tables to scale point-query-only based workloads to modern SSDs, but lacks first-class support for range operations.
-KVell (Lepers et al., 2019), later extended to KVell+ (Lepers et al., 2020), achieves high performance on modern SSDs using a shared-nothing design based on independent sharded indices maintained on separate threads.
+SplinterDB (conwaySplinterDBClosingBandwidth2020) and Mapped SplinterDB (conwaySplinterDBMapletsImproving2023) introduce innovations in the design of external memory dictionaries, extending Bϵ-Trees to STBϵ-Trees by applying ideas from size-tiered LSM-trees to achieve high update efficiency without excessively penalizing point queries.
+FASTER (chandramouliFASTERConcurrentKeyValue2018; chandramouliFASTEREmbeddedConcurrent2018) combines log-structured storage with in-memory concurrent hash tables to scale point-query-only based workloads to modern SSDs, but lacks first-class support for range operations.
+KVell (lepersKVellDesignImplementation2019a), later extended to KVell+ (lepersKVellSnapshotIsolation2020), achieves high performance on modern SSDs using a shared-nothing design based on independent sharded indices maintained on separate threads.
 KVell only maintains key-based indices in memory (not in storage), trading scalability and recovery time for lower space-/write-amplification.